### v1

# Contrastive explainable clustering with differential privacy

Abstract
This paper presents a novel approach in Explainable AI (XAI), integrating contrastive explanations with differential privacy in clustering methods. For several basic clustering problems, including $k$-median and $k$-means, we give efficient differential private contrastive explanations that achieve essentially the same explanations as those that non-private clustering explanations can obtain. We define contrastive explanations as the utility difference between the original clustering utility and utility from clustering with a specifically fixed centroid. In each contrastive scenario, we designate a specific data point as the fixed centroid position, enabling us to measure the impact of this constraint on clustering utility under differential privacy. Extensive experiments across various datasets show our method’s effectiveness in providing meaningful explanations without significantly compromising data privacy or clustering utility. This underscores our contribution to privacy-aware machine learning, demonstrating the feasibility of achieving a balance between privacy and utility in the explanation of clustering tasks.
1 Introduction
Different notions of clustering are fundamental primitives in several areas, including machine learning, data science, and operations research Reddy (2018). Clustering objectives (e.g., $k$-median and $k$-means), crucial in facility location planning, have been notably applied in pandemic response strategies, such as determining optimal sites for COVID-19 vaccination and testing centers  Bertsimas et al. (2022); Mehrab et al. (2022). The distance to a cluster center becomes a significant factor in these scenarios, particularly highlighted during the COVID-19 pandemic, where accessible vaccine locations played a pivotal role  Rader et al. (2022). This context often leads to individuals, especially those distant from a cluster center, seeking explanations for their specific assignment or location in the clustering plan. For example, consider a resident questioning why a COVID-19 testing center isn’t located next to their home. Such queries reflect the need for explainability in clustering decisions, where factors like population density, resource allocation, and public health considerations impact site selection. The resident’s inquiry underscores the importance of transparent methodologies in clustering that can provide understandable explanations for facility placements, especially in critical health scenarios.
Such questions fall within the area of Explainable AI, which is a rapidly growing and vast area of research Finkelstein et al. (2022); Sreedharan et al. (2020); Miller (2019); Boggess et al. (2023). We focus on post-hoc explanations, especially contrastive explanations, e.g., Miller (2019), which address “why P instead of Q?” questions. Such methods have been used in many applications, including multi-agent systems, reinforcement learning, and contrastive analysis, e.g., Boggess et al. (2023); Sreedharan et al. (2021); Madumal et al. (2020); Sreedharan et al. (2020). In the context of the clustering problems mentioned above, a possible explanation for a resident is the increase in the cost of the clustering if a center is placed close to the resident, compared to the cost of an optimal solution, which is relatively easy to compute. This increased cost reflects a trade-off, where optimizing for one resident’s convenience could result in worse centroid positions for other residents, highlighting the complex decision-making process in clustering for fair resource distribution.
Data privacy is a crucial concern across various fields, and Differential Privacy (DP) is one of the most widely used and rigorous models for privacy Dwork and Roth (2014). We focus on the setting where the set of data points $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ are private; for instance, in the vaccine center deployment problem, each $x_{i}$ corresponds to an individual who is seeking vaccines, and would like to be private. There has been a lot of work on the design of differentially private solutions to clustering problems like $k$-median and $k$-means under such a privacy model Ghazi et al. (2020); Gupta et al. (2010); Balcan et al. (2017); Stemmer and Kaplan (2018). These methods guarantee privacy of the data points $\mathbf{X}$ and output a set of centers $c_{1},\ldots,c_{k}$ (which are public) such that its cost, denoted by $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})$ is within multiplicative and additive factors $w$ and $t$ of the optimal solution $OPT_{\mathbf{X}}$, i.e., $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})\leq w\cdot OPT_{\mathbf{X}}+t$.
While there has been significant progress in various domains of differential privacy, the intersection of explainability and differential privacy still needs to be explored. In clustering problems, a potential private contrastive explanation given to agent $i$ is to provide the $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$, where $S_{\epsilon}$ is a set of centers generated by private algorithm (with privacy budget $\epsilon$), and $\text{cost}(S^{(i)}_{\epsilon})$ is the cost of the clustering solution while fixing one center at agent $i$’s location. However, giving such a private contrastive explanation to each agent $i$ naively using a $(w,t)$-approximate private clustering would lead to $(w,t\sqrt{n}/\epsilon)$-approximation, even if we use advanced composition techniques for privacy. The central question of our research: is it possible to offer each user a private contrastive explanation with a limited overall privacy budget?
Our contributions.
1. We introduce the PrivEC problem, designed to formalize private contrastive explanations to all agents in clustering using $k$-median and $k$-means objectives.
2. We present an $\epsilon$-DP mechanism, PrivateExplanations, which provides a contrastive explanation to each agent while ensuring the same utility bounds as private clustering in Euclidean spaces Ghazi et al. (2020). We use the private coreset technique of Ghazi et al. (2020), which is an intermediate private data structure that preserves similar clustering costs as the original data. The main technical challenge is to derive rigorous bounds on the approximation factors for all the contrastive explanations.
3. We evaluate our methods empirically on realistic datasets used in the problems of vaccine center deployment in Virginia Mehrab et al. (2022). Our results show that the privacy-utility trade-offs are similar to those in private clustering, and the errors are quite low even for reasonable privacy budgets, demonstrating the effectiveness of our methods.
Our research stands out by seamlessly integrating differential privacy into contrastive explanations, maintaining the quality of explanations even under privacy constraints. This work bridges the gap between privacy and explainability, marking a significant advancement in privacy-aware machine learning. Due to space limitations, we present many technical details and experimental evaluations in the Appendix.
2 Related Work
Our work considers differential privacy for explainable AI in general (XAI) and Multi-agent explanations (XMASE) in particular, focusing on post-hoc contrastive explanations for clustering. We summarize some of the work directly related to our paper; additional discussion is presented in the Appendix, due to space limitations.
Extensive experiments presented in Saifullah et al. (2022) demonstrate non-negligible changes in explanations of black-box ML models through the introduction of privacy.
Nguyen et al. (2023) considers feature-based explanations (e.g., SHAP) that can expose the top important features that a black-box model focuses on. To prevent such expose they introduced a new concept of achieving local differential privacy (LDP) in the explanations, and from that, they established a defense, called XRAND, against such attacks. They showed that their mechanism restricts the information that the adversary can learn about the top important features while maintaining the faithfulness of the explanations.
Goethals et al. (2022) study the security of contrastive explanations, and introduce the concept of the “explanation linkage attack”, a potential vulnerability that arises when employing strategies to derive contrastive explanations. To address this concern, they put forth the notion of k-anonymous contrastive explanations. As the degree of privacy constraints increases, a discernible trade-off comes into play: the quality of explanations and, consequently, transparency are compromised.
Closer to our application is the work of Georgara et al. (2022), which investigates the privacy aspects of contrastive explanations in the context of team formation. They present a comprehensive framework that integrates team formation solutions with their corresponding explanations, while also addressing potential privacy concerns associated with these explanations. Additional evaluations are needed to determine the privacy of such heuristic-based methods.
There has been a lot of work on private clustering and facility location, starting with  Gupta et al. (2010), which was followed by a lot of work on other clustering problems in different privacy models, e.g., Huang and Liu (2018); Stemmer (2020); Stemmer and Kaplan (2018); Nissim and Stemmer (2018); Feldman et al. (2017). Gupta et al. (2010) demonstrated that the additive error bound for points in a metric space involves an $O(\Delta k^{2}\log{n}/\epsilon)$ ) term, where $\Delta$ is the space’s diameter. Consequently, all subsequent work, including ours, assumes points are restricted to a unit ball.
We note that our problem has not been considered in any prior work in the XAI or differential privacy literature. The formulation we study here will likely be useful for other problems requiring private contrastive explanations.
3 Preliminaries
Let $\mathbf{X}\subset\mathbb{R}^{d}$ denote a dataset consisting of $d$-dimensional points (referred to sometimes as agents). We consider the notion of $(k,p)$-clustering, as defined by Ghazi et al. (2020).
Definition 1.
($(k,p)$-Clustering Ghazi et al. (2020)). Given $k\in\mathbb{N}$, and a multiset $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ of points in the unit ball, a $(k,p)$-clustering is a set of $k$ centers $\{c_{1},\ldots,c_{k}\}$ minimizing $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})=\sum_{i\in[n]}\min_{j\in[k]}\|x_{i}-c_{j}\|^{p}$.
For $p=1$ and $p=2$, this corresponds to the $k$-median and $k$-means objectives, respectively. We drop the subscript $\mathbf{X}$ and superscript $p$, when it is clear from the context, and refer to the cost of a feasible clustering solution $S$ by $\text{cost}(S)$.
Definition 2.
($(w,t)$-approximation). Given $k\in\mathbb{N}$, and a multiset $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ of points in the unit ball, let $OPT^{p,k}_{\mathbf{X}}=\min_{c_{1},\ldots,c_{k}\in\mathbb{R}^{d}}\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})$ denote the cost of an optimal $(k,p)$-clustering. We say $c_{1},\ldots,c_{k}$ is a $(w,t)$– approximation to a $(k,p)$-optimal clustering if $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})\leq w\cdot OPT^{p,k}_{\mathbf{X}}+t$.
A coreset (of some original set) is a set of points that, given any $k$ centers, the cost of clustering of the original set is “roughly” the same as that of the coreset Ghazi et al. (2020).
Definition 3.
For $\gamma,t>0,p\geq 1,k,d\in\mathbb{N}$, a set $X^{\prime}$ is a $(p,k,\gamma,t)$-coreset of $X\subseteq\mathbb{R}^{d}$ if for every $C=\{c_{1},\ldots,c_{k}\}\in R^{d}$, we have $(1-\gamma)cost_{X}^{p}(C)-t\leq cost_{X^{\prime}}(C)\leq(1+\gamma)cost_{X}^{p}(C)+t$.
Privacy model. We use the notion of differential privacy (DP), introduced in Dwork and Roth (2014), which is a widely accepted formalization of privacy. A mechanism is DP if its output doesn’t differ too much on “neighboring” datasets; this is formalized below.
Definition 4.
$\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-differentially private if for any neighboring datasets $X\sim X^{\prime}\in\mathcal{X}$ and $S\subseteq\mathcal{Y}$,
$\Pr[\mathcal{M}(X)\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(X^{\prime})\in S]+\delta.$
If $\delta=0$, we say $\mathcal{M}$ is $\epsilon$-differentially private.
We operate under the assumption that the data points in $\mathbf{X}$ represent the private information of users or agents. We say $\mathbf{X},\mathbf{X}^{\prime}$ are neighboring (denoted by $\mathbf{X}\sim\mathbf{X}^{\prime}$) if they differ in one data point. When a value is disclosed to an individual agent $i$, it is imperative to treat the remaining clients in $\mathbf{X}-\{i\}$ as private entities. To address this necessity, we adopt the following privacy model:
Definition 5.
A mechanism $\mathcal{M}$ is $\epsilon$-$i$-exclusive DP if, $\forall X,X^{\prime}:X\setminus\{i\}\sim X^{\prime}\setminus\{i\}$, and for all $S\subseteq Range(\mathcal{M})$:
$\Pr[\mathcal{M}(X)\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(X^{\prime})\in S].$
Note that if $\mathcal{M}$ is $\epsilon$-DP, it is $\epsilon$-$i$-exclusive DP for any $i$.
Relation to Joint-DP model. Kearns et al. (2014) defines Joint-DP–a privacy model that “for each of agent $i$, knowledge of the $n-1$ agents does not reveal much about agent $i$”. It contrasts with our privacy model, where each explanation to agent $i$ uses direct and private knowledge of agent $i$ (that private knowledge will be returned to only agent $i$) and such knowledge does not reveal much about $n-1$ other agents.
In clustering scenarios such as k-means or k-median, it is beneficial for an individual data point to be positioned near the cluster center, as this reduces its distance. This concept has practical applications in fields such as urban planning and public health. For example, clustering algorithms like $k$-means and $k$-median have been utilized to optimize the placement of mobile vaccine centers, as noted in  Mehrab et al. (2022). However, once the cluster centers are calculated, residents might wonder why a vaccine clinic has not been positioned nearer to their homes. Our objective is to provide explanations for each data point that is not near a cluster center. Following the approach suggested by  Miller (2019), one way to provide an explanation is to re-run the clustering algorithm, this time setting the resident’s location as a fixed centroid. This allows us to explain the changes in clustering costs, where an increase could describe why certain placements are less optimal.
The focus of this paper is twofold: residents not only need explanations but also wish to maintain their privacy, preferring not to disclose their precise locations/data publicly. Traditional private clustering methods struggle to offer these private explanations because they accumulate the privacy budget by continuously applying private clustering to answer numerous inquiries, ultimately resulting in explanations that are not private. Thus, we have developed a framework that delivers the necessary explanations to residents while protecting their confidentiality.
In essence, our framework is data-agnostic and capable of offering contrastive explanations for any private data within clustering algorithms, effectively balancing the need for privacy with the demand for transparency in decision-making.
Definition 6.
Private and Explainable Clustering problem (PrivEC) Given an instance $\mathbf{X}\in\mathcal{X}^{n}$, clustering parameters $k,p$, and a subset of data points: $V_{s}\subset\mathbf{X}$, the goal is to output:
Private: An $\epsilon$-DP clustering solution $S_{\epsilon}$ (available to all)
Explainable: For each $i\in V_{s}$, output $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$, which is $\epsilon$-$i$-exclusive DP.
$S^{(i)}_{\epsilon}$ is a private solution computed by the clustering algorithm while fixing one centroid to the position of agent $i$.
We assume that $S^{(i)}_{\epsilon}$ is not revealed to any agent, but $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$ is released to agent $i$ as contrastive explanation.
4 peCluster Mechanism
The PrivateClustering algorithm Ghazi et al. (2020) includes dimension reduction, forming a private coreset, applying an approximation clustering algorithm on the coreset, and then reversing the dimension reduction, ultimately yielding an $\epsilon$-DP clustering solution.
The PrivateExplanations algorithm receives the original private clustering cost from PrivateClustering and then leverages the PrivateClustering framework but introduces a fixed-centroid clustering algorithm on the coreset instead of regular approximation clustering algorithm, followed by a dimension reverse. Ultimately, after calculating the fixed centroid clustering algorithm and determining its cost in the original dimension, we subtract the original private clustering cost obtained from the PrivateClustering algorithm, outputting $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$.
To elaborate further, Private Clustering first reduces the dataset’s dimensionality using DimReduction (Algorithm 1) to enable efficient differentially private coreset construction. It then uses a DP Exponential Mechanism to build a PrivateCoreset (Algorithm 7 of Ghazi et al. (2020)), which is clustered using any $w^{\prime\prime}$-approximate, not necessarily private clustering algorithm. The low-dimensional clustering cost is scaled to obtain the original high-dimensional cost, and the centroids are recovered in the original space. Multiplying the low-dimensional clustering cost with $(\log(n/\beta)/0.01)^{p/2}$ gives the clustering cost of the original high-dimensional instance Makarychev et al. (2019). By Running DimReverse using FindCenter (Algorithm 10 of Ghazi et al. (2020)) we can find the centroids in the original space. The output is the private centroids, their cost, the private coreset, and the low-dimensional dataset. We employed DimReduction since PrivateCoreset is an exponential algorithm. This approach enables the transformation of dimensionality reduction into an exponential process, effectively converting the algorithm’s time complexity from exponential to polynomial with respect to the input size.
The following summarizes Algo. 3 - PrivateClustering:
Input: Agents’ location ($X$), original and reduced dimension (d, d’). Output: Private clustering: centers and cost.
  1. 1.
    Dimension reduction - Algorithm 1 DimReduction
  2. 2.
    Private coreset - Algorithm. PrivateCoreset
  3. 3.
    Clustering on the coreset - Algorithm NonPrivateApprox
  4. 4.
    Dimension reverse -  2 DimReverse
Algorithm 1 DimReduction
Input: $(x_{1},x_{2},\ldots,x_{n}),d,d^{\prime},\beta$
Output: $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$ low-dimensional space dataset.
1:  $\Lambda=\sqrt{\frac{0.01d}{\log(n/\beta)d^{\prime}}}$
2:  for $i\in\{1,..,n\}$ do
3:     $\tilde{x}_{i}\leftarrow\Pi_{\mathcal{S}}(x_{i})$
4:     if $\|\tilde{x}_{i}\|\leq 1/\Lambda$ then
5:        $x^{\prime}_{i}=\Lambda\tilde{x}_{i}$
6:     else
7:        $x^{\prime}_{i}=0$
8:     end if
9:  end for
10:  return $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
Algorithm 2 DimReverse
Input: $(c^{\prime}_{1},\ldots,c^{\prime}_{k}),(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
Output: $(c_{1},\ldots,c_{k})$ Private Centroids in high dimension
1:  $\mathbf{X}_{1},\ldots,\mathbf{X}_{k}\leftarrow\text{ the partition induced by }$ $(c^{\prime}_{1},\ldots,c^{\prime}_{k})$ on $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
2:  for $j\in\{1,\ldots,k\}$ do
3:     $c_{j}\leftarrow\textsc{FindCenter}^{\epsilon/2}(\mathbf{X}_{j})$
4:  end for
5:  return $(c_{1},\ldots,c_{k})$
Algorithm 3 PrivateClustering
Input: $(x_{1},x_{2},\ldots,x_{n}),d,d^{\prime},k,\epsilon,\lambda,p,\alpha,\beta$
 Output: $(\epsilon,\delta)$-differentially private centers $c_{1,..,k}$ with cost $S_{\epsilon}$, $Y$ Private Coreset and $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$ low-dimensional space dataset.
1:  $\zeta=0.01\left(\frac{\alpha}{10\lambda_{p,\alpha/2}}\right)^{p/2}$
2:  $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$leftarrow $DimReduction((x_{1},\ldots,x_{n}),d,d^{\prime},\beta)$
3:  $Y\leftarrow\textsc{PrivateCoreset}^{\epsilon/2}(x^{\prime}_{1},\ldots,x^{\prime}_{n},\zeta)$
4:  $((c^{\prime}_{1},\dots,c^{\prime}_{k}),cost(S^{\prime}_{\epsilon}))\leftarrow NonPrivateApprox(Y,k)$
5:  $cost(S_{\epsilon})=\left(\frac{\log(n/\beta)}{0.01}\right)^{p/2}cost(S^{\prime}_{\epsilon})$
6:  $c\leftarrow\textsc{DimReverse}^{\epsilon/2}((c^{\prime}_{1},\dots,c^{\prime}_{k}),(x^{\prime}_{1},\dots,x^{\prime}_{n}))$
7:  return $c=\{c_{1},\ldots,c_{k}\},cost(S_{\epsilon})$, $Y$, $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
Algorithm 4 PrivateExplanations
Input: $Y,\text{cost}(S_{\epsilon}),V_{s},k,\epsilon,p,n,\beta$
 Output: {$cost(S_{\epsilon})-cost(S^{(i)}_{\epsilon}):i\in V_{s}\}$ DP explanations for each agent (i)
1:  for $i\in V_{s}$ do
2:     $cost(S^{(i)}_{\epsilon})\leftarrow\left(\frac{\log(n/\beta)}{0.01}\right)^{\frac{p}{2}}NonPrivateApproxFC(Y,k,x^{\prime}_{i})$
3:  end for
4:  return $\{cost(S_{\epsilon})-cost(S^{(i)}_{\epsilon}):i\in V_{s}\}$
Private contrastive explanation:
The following summarizes Algorithm 4 - PrivateExplanations:
Input: Private coreset ($Y$), Private clustering cost ($cost(S_{\epsilon})$), Agents looking for contrastive explanation ($V_{s}$).
Output: Explanation for each agent in $V_{s}$ ( $\text{cost}(S_{\epsilon})-cost(S^{(i)}_{\epsilon}$)).
  1. 1.
    Fixed-centroid clustering on Coreset - NonPrivateApproxFC (specifically for $k$-means and $k$-median in Section 4.2 and Section 4.1 respectively)
  2. 2.
    Calculate clustering cost by multiplication - Algorithm 4, line 2.
  3. 3.
    Return utility decrease as an explanation - Algorithm 4, line 4.
After the completion of PrivateClustering, which yields private centroids, specific users within $V_{s}\subset\mathbf{X}$, may request an explanation regarding the absence of a facility at their particular location. We want to provide each agent with the overall cost overhead that a facility constraint at his specific location will yield. To address these queries, we integrate our peCluster mechanism. This algorithm takes the previously computed coreset and employs our NonPrivateApproxFC(tailor-suited non-private algorithm we designed and explained in the following sections) $w^{\prime}$-approximate algorithm to output the cost of $(k,p)$-clustering with a fixed centroid.
The output of our algorithm provides the difference cost of the clustering, denoted as cost($S^{(i)}_{\epsilon}$) - cost($S_{\epsilon}$). This output effectively captures the contrast between the overall clustering cost and the specific clustering cost when an individual agent ($i$) inquires about the absence of a facility at their location, aligning precisely with the key objectives and functions outlined earlier in the abstract.
The privacy analysis, as demonstrated in Theorem 1, establishes the privacy guarantees of PrivateClustering and peCluster. The output $Y$ is $\epsilon/2$-differentially private as an output of $\epsilon/2$-DP algorithm. Consequently, both $(c^{\prime}_{1},\ldots,c^{\prime}_{k})$ and $cost(S_{\epsilon})$ maintain $\epsilon/2$-DP status, under the Post-Processing property.
Applying $\textsc{DimReverse}^{\epsilon/2}$ to find the centers in the original space, $c=\{c_{1},\ldots,c_{k}\}$ is $\epsilon$-DP by Composition theorem. For each $i$, cost($S^{(i)}_{\epsilon}$) is produced by Post-Processing of $Y$ with only $x^{\prime}_{i}$, hence $cost(S^{(i)}_{\epsilon}$) satisfies $\epsilon$-$i$-exclusive-DP.
Theorem 1.
(Full proof in Theorem 5.) The solution $(c_{1},\ldots,c_{k})$ and $cost(S_{\epsilon})$ released by Algorithm 3 are $\epsilon$-DP. For all clients $i$, the value $S^{(i)}_{\epsilon}$ released by Algorithm 4 is $\epsilon$-$x_{i}$-exclusive DP.
A critical component of our algorithm involves a detailed utility analysis. In the following sections, we will present rigorous upper bounds and specific constraints for k-means and k-median.
Utility Analyses. PrivateCoreset is setup with parameter $\zeta$ (other than the privacy budget $\epsilon/2$), which depends on $\alpha$ (Line 1, Algorithm 3). It leads to the cost of every clustering on $Y$ (the coreset) is off from its cost on $X^{\prime}$ by a multiplicative factor of $(1+0.1\alpha)$ plus an additive factor of $\overset{\sim}{O}(\text{polylog}(n/\beta)/\epsilon)$. Then, by applying the Dimensional Reduction lemma (Appendix section B), which states that the cost of a specific clustering on $X^{\prime}$ ($d^{\prime}$-dimensional space) is under some constant factor of the same clustering on $X$ ($d$-dimensional space), we can bound the $cost({S_{\epsilon}}^{(i)})$ by its optimal clustering $OPT_{i}$.
Lemma 1.
(Full proof in Lemma 7) With probability at least $1-\beta$, $S_{\epsilon}$ (clustering cost) is a $(w,t)$-approximation of $OPT$, where [^1] We use the notation $\mathcal{O}_{p,\alpha}$ to explicitly ignore factors of $p,\alpha$:
$w$ $=w^{\prime\prime}(1+\alpha),$ (1)
$t$ $=w^{\prime\prime}\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.\text{polylog}(n/\beta)/\epsilon\right)$ (2)
We note that $\Lambda$ is chosen carefully by composing (1) the factor of $(d/d^{\prime})^{1/2}$, which will scale to $(d/d^{\prime})^{p/2}$ due to .$\|.\|_{p}$ operator and cancel out the scaling factor $(d^{\prime}/d)^{p/2}$ occurs when mapping the cost of clustering from the lower dimensional space to the original space and (2) the factor of $0.01/\log(n/\beta)$ that guarantees each $\|x^{\prime}_{i}\|<1/\Lambda$ with probability at least $1-0.1\beta/n$ (and for all $i$ simultaneously with probability $1-0.1\beta$).
Theorem 2.
(Full proof in Theorem 6) Fix an $i$. With probability at least $1-\beta$, $S^{(i)}_{\epsilon}$ released by Algorithm 4 is a $(w,t)$-approximation of $OPT_{i}$-the optimal $(k,p)$-clustering cost with a center fixed at position $x_{i}$, in which:
$w$ $=w^{\prime}(1+\alpha)$ (3)
$t$ $=w^{\prime}\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.\text{polylog}(n/\beta)/\epsilon\right)$ (4)
Lemma 2.
(Full proof in Lemma 8) Fix an $i$. If $OPT_{i}\geq w^{\prime\prime}(1+\alpha)OPT+t^{(i)}$, then with probability at least $1-2\beta$, ${S_{\epsilon}}$ and ${S_{\epsilon}}^{(i)}$ released by Algorithm 3 and 4 satisfies that ${S_{\epsilon}}^{(i)}>{S_{\epsilon}}$.
Running Time Analysis. Assume that polynomial-time algorithms exist for $(k,p)$-clustering and for $(k,p)$-clustering with a fixed center. Setting $d^{\prime}=O(p^{4}\log(k/\beta))$ (that satisfies Dimension-Reduction-Lemma(Appendix section B)), algorithm PrivateCoreset takes $O(2^{O(d^{\prime})}\text{poly}(n))$ time (Lemma 42 of Ghazi et al. (2020)), which is equivalent to $O((k/\beta)^{O_{p,\alpha}(1)}\text{poly}(n))$. FindCenter takes $O(\text{poly}(np))$ time (Corollary 54 of Ghazi et al. (2020)) and is being called $k$ times in total. Finally, we run $1$ instance of $(k,p)$-clustering and $|V_{s}|$ instances of $(k,p)$-clustering with a fixed center. It follows that the total running time of our algorithm is $O((k/\beta)^{O_{p,\alpha}(1)}\text{poly}(nd))$. This means the computational complexity of our algorithm is polynomial in relation to the size of the input data.
Tight Approximation Ratios. Corollary 1 and 2 will wrap up this section by presenting a specific, tight approximation ratio($w^{\prime}$) achieved after applying NonPrivateApproxFC. These corollaries will detail and confirm the algorithm’s effectiveness in achieving these approximation ratios.
4.1 NonPrivateApproxFC for $k$-median
As previously mentioned, we have specifically developed a non-private fixed centroid clustering algorithm, referred to as NonPrivateApproxFC. Our algorithm is proved to be an 8-approximation algorithm. We obtain our results by adapting Charikar et al. (1999) to work with a fixed centroid (Moving forward in our discussion, we will refer to the fixed centroid as $C$). To grasp how we adapted the algorithm to suit our needs, it’s essential to understand the definitions used in  Charikar et al. (1999) as they introduce demand $d_{j}$ for each location $j\in N$ as a weight reflecting the location’s importance. $N$ represents list of agents ${1\ldots n}$. For the conventional k-median problem, each $d_{j}$ is initially set to 1 for all $j\in N$. The term $c_{ij}$ represents the cost of assigning any $i$ to $j$ and $x_{ij}$ represents if location $j$ is assigned to center $i$.
We first assume that the fixed center is one of the input data points $N$ to formulate the $k$-median problem with a fixed center as the solution of the integer program (in the Appendix). We can then relax the IP condition to get the following LP:
minimize $\sum_{i,j\in N}d_{j}c_{ij}x_{ij}$ (5)
[…]
$x_{ij}$ $\leq y_{i}\text{ for each $i,j\in N$}$ (7)
$x_{ij},y_{i}$ $\geq 0\text{ for each $i,j\in N$}$ (8)
$y_{C},x_{CC}$ $\geq 1\text{ for a fixed $C\in N$}$ (9)
Let $(\bar{x},\bar{y})$ be a feasible solution of the LP relaxation and let $\bar{C}_{j}=\sum_{i\in N}c_{ij}\bar{x}_{ij}$ for each $j\in N$ as the total (fractional) cost of client $j$.
The first step. We group nearby locations by their demands without increasing the cost of a feasible solution $(\bar{x},\bar{y})$, such that locations with positive demands are relatively far from each other. By re-indexing, we get $\bar{C}_{C}\leq\bar{C}_{1}\leq\bar{C}_{2}\leq\ldots\bar{C}_{n}$.
We will show that it’s always possible to position $\bar{C}_{C}$ as the first element of the list, i.e., $\bar{C}_{C}$ is equal to the minimum value of all $\bar{C}_{j}$. Recall that: $\bar{C}_{C}=\sum_{i\in N}c_{iC}\bar{x}_{iC}=\sum_{i\in N,i\neq C}c_{iC}\bar{x}_{iC}+c_{CC}\bar{x}_{CC}=0,$ since we know that $\sum_{i\in N}x_{iC}=1,x_{CC}\geq 1$ and $c_{CC}=0$.
The remaining work of the first step follows Charikar et al. (1999). We first set the modified demands ${d_{j}}^{\prime}\leftarrow d_{j}$. For $j\in N$, moving all demand of location $j$ to a location $i<j$ s.t. $d^{\prime}_{i}>0$ and $c_{ij}\leq 4\bar{C}_{j}$, i.e., transferring all $j$’s demand to a nearby location with existing positive demand. Demand shift occurs as follows: $d^{\prime}_{i}\leftarrow d^{\prime}_{i}+d^{\prime}_{j},d^{\prime}_{j}\leftarrow 0$. Since we initialize $d^{\prime}_{C}=d_{C}=1$, and we never move its demands away, it follows that $d^{\prime}_{C}>0$.
Let $N^{\prime}$ be the set of locations with positive demands $N^{\prime}=\{j\in N,d^{\prime}_{j}>0\}$. A feasible solution to the original demands is also a feasible solution to the modified demands. By Lemma 11, for any pair $i,j\in N^{\prime}$: $c_{ij}>4\max(\bar{C}_{i},\bar{C}_{j})$.
Lemma 3.
The cost of the fractional $(\bar{x},\bar{y})$ for the input with modified demands is at most its cost for the original input.
Proof.
The cost of the LP $\bar{C}_{LP}=\sum_{j\in N}d_{j}\bar{C}_{j}$ and $\bar{C}^{\prime}_{LP}=\sum_{j\in N}d^{\prime}_{j}\bar{C}_{j}$. Since we move the demands from $\bar{C}_{j}$ to a location $i$ with lower cost $\bar{C}_{i}\leq\bar{C}_{j}$ the contribution of such moved demands in $\bar{C}^{\prime}$ is less than its contribution in $\bar{C}$, it follows that $\bar{C}^{\prime}_{LP}\leq\bar{C}_{LP}$. ∎
The second step. We analyze the problem with modified demands $d^{\prime}$. We will group fractional centers from the solution $(\bar{x},\bar{y})$ to create a new solution $(x^{\prime},y^{\prime})$ with cost at most $2\bar{C}_{LP}$ such that $y^{\prime}_{i}=0$ for each $i\notin N^{\prime}$ and $y^{\prime}_{i}\geq 1/2$ for each $i\in N^{\prime}$. We also ensure that $y^{\prime}_{C}\geq 1/2$ in this step, i.e., $C$ will be a fractional center after this.
The next lemma states that for each $j\in N$, there are at least $1/2$ of the total fractional centers within the distance of $2\bar{C}_{j}$.
Lemma 4.
For any $1/2$-restricted solution $(x^{\prime},y^{\prime})$ there exists a $\{1/2,1\}$-integral solution with no greater cost.
Proof.
The cost of the $\frac{1}{2}$-restricted solution (by Lemma 7 of Charikar et al. (1999)) is:
$C^{\prime}_{LP}$ $=\sum_{j\in N^{\prime}}d^{\prime}_{j}c_{s(j)j}-\sum_{j\in N^{\prime}}d^{\prime}_{j}c_{s(j)j}y^{\prime}_{j},$ (10)
Let $s(j)$ be $j$’s closest neighbor location in $N^{\prime}$, the first term above is independent of $y^{\prime}$ and the minimum value of $y^{\prime}_{j}$ is $1/2$. We now construct a $\{1/2,1\}$-integral solution $(\hat{x},\hat{y})$ with no greater cost. Sort the location $j\in N^{\prime},j\neq C$ in the decreasing order of the weight $d^{\prime}_{j}c_{c_{s}(j)j}$ and put $C$ to the first of the sequence, set $\hat{y}_{j}=1$ for the first $2k-n^{\prime}$ locations and $\hat{y}_{j}=1/2$ for the rest. By doing that, we minimize the cost by assigning heaviest weights $d^{\prime}_{j}c_{c_{s}(j)j}$ to the maximum multiplier (i.e., $1$) while assigning lightest weights $d^{\prime}_{j}c_{c_{s}(j)j}$ to the minimum multiplier (i.e., $1/2$) for each $j\in N^{\prime},j\neq C$. Any feasible $1/2$-restricted solution must have $y^{\prime}_{C}=1$ to satisfy the constraint of $C$ so that the contribution of $\hat{y}_{C}$ is the same as its of $y^{\prime}_{C}$. It follows that the cost of $(\hat{x},\hat{y})$ is no more than the cost of $(x^{\prime},y^{\prime})$. ∎
The third step. This step is similar to the part of Step 3 of Charikar et al. (1999) that converts a $\{1/2,1\}$-integral solution to an integral solution with the cost increases at most by $2$. We note that there are two types of center $\hat{y}_{j}=1/2$ and $\hat{y}=1$, hence there are two different processes. All centers $j$ with $\hat{y}_{j}=1$ are kept while more than half of centers $j$ with $\hat{y}_{j}=1/2$ are removed. Since we show that $\hat{y}_{C}=1$ in the previous step, $C$ is always chosen by this step and hence guarantees the constraint of $C$.
Theorem 3.
For the metric k-median problem, the algorithm above outputs an $8$-approximation solution.
Proof.
It is obvious that the optimal of the LP relaxation is the lower bound of the optimal of the integer program. While constructing an integer solution for the LP relaxation with the modified demands, Lemma 13 incurs a cost of $2\bar{C}_{LP}$ and the third step multiplies the cost by a factor of $2$, making the cost of the solution (to the LP) at most $4\bar{C}_{LP}$. Transforming the integer solution of the modidied demands to a solution of the original input adds an additive cost of $4\bar{C}_{LP}$by Lemma 12 and the Theorem follows. ∎
Corollary 1.
Running peCluster with NonPrivateApproxFC be the above K-median algorithm, with probability at least $1-\beta$, $S^{(i)}_{\epsilon}$ is a $(w,t)$-approximation of $OPT_{i}$–the optimal K-median with a center fixed at position $x_{i}$, in which:
$w$ $=8(1+\alpha)$
$t$ $=8\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.polylog(n/\beta)/\epsilon\right).$
4.2 NonPrivateApproxFC for $k$-means
In this section, we describe an algorithm that is a $25$-approximation of the k-means problem with a fixed center. We adapt the work by Kanungo et al. (2002) by adding a fixed center constraint to the single-swap heuristic algorithm. Similar to Kanungo et al. (2002), we need to assume that we are given a discrete set of candidate centers $C$ from which we choose $k$ centers. The optimality is defined in the space of all feasible solutions in $C$, i.e., over all subsets of size $k$ of $C$. We then present how to remove this assumption, with the cost of a small constant additive factor, in the Appendix.
Definition 7.
Let $O=(O_{1},O_{2},\ldots,O_{k})$ be the optimal clustering with $O_{1}$ be the cluster with the fixed center $\sigma$. A set $C\subset\mathbb{R}^{d}$ is a $\gamma$-approximate candidate center set if there exists $\sigma\in\{c_{1},c_{2},\ldots,c_{k}\}\subseteq C$, such that:
$cost(c_{1},c_{2},\ldots,c_{k})$ $\leq(1+\gamma)cost(O),$ (11)
Given $u,v\in\mathbb{R}^{d}$, let $\Delta(u,v)$ denote the squared Euclidean distance between $u,v:\Delta(u,v)=dist^{2}(u,v)$. Let $S\subset\mathbb{R}^{d}$, $\Delta(S,v)=\sum_{u\in S}\Delta(u,v)$. Let $P\subset\mathbb{R}^{d}$, $\Delta_{P}(S)=\sum_{q\in P}\Delta(q,s_{q})$, where $s_{q}$ is the closest point of $S$ to $q$. We drop $P$ when the context is clear.
Let $\sigma$ be the fixed center that must be in the output. Let $C$ be the set of candidate centers, that $\sigma\in C$. We define stability in the context of $k$-means with a fixed center $\sigma$ as follows. We note that it differs from the definition of Kanungo et al. (2002) such that we never swap out the fixed center $\sigma$:
Definition 8.
A set $S$ of k centers that contains the fixed center $\sigma$ is called $1$-stable if: $\Delta\big{(}S\setminus\{s\}\cup\{o\}\big{)}\geq\Delta(S),$ for all $s\in S\setminus\{\sigma\}$, $o\in O\setminus\{\sigma\}$.
Algorithm. We initialize $S^{(0)}$ as a set of $k$ centers form $C$ that $\sigma\in S^{(0)}$. For each set $S^{(i)}$, we perform the swapping iteration:
  • Select one center $s\in S^{(i)}\setminus\sigma$
  • Select one replaced center $s^{\prime}\in C\setminus S^{(i)}$
  • Let $S^{\prime}=S^{(i)}\setminus s\cup s^{\prime}$
  • If $S^{\prime}$ reduces the distortion, $S^{(i+1)}=S^{\prime}$. Else, $S^{(i+1)}=S^{(i)}$
We repeat the swapping iteration until $S=S^{(m)}$, i.e., after $m$ iterations, is a $1$-stable. Theorem 4 states the utility of an arbitrary $1$-stable set, which is also the utility of our algorithm since it always outputs an $1$-stable set. We note that if $C$ is created with some errors $\gamma$ to the actual optimal centroids, the utility bound of our algorithm is increased by the factor $\Theta(\gamma)$, i.e., ours is a $(25+\Theta(\gamma))$-approximation to the actual optimal centroids.
Theorem 4.
If $S$ is an $1$-stable $k$-element set of centers, then $\Delta(S)\leq 25\Delta(O)$. Furthermore, if $C$ is a $\frac{\gamma}{25}$-approximate candidate center set, then $S$ is a $(25+\gamma)$-approximate of the actual optimal centroids in the Euclidean space.
Kanungo et al. (2002)’s analysis depends on the fact that all centers in the optimal solution are centroidal, hence they can apply centroidal-Lemma in the Appendix. Our analysis must take into account that $\sigma\in O$ is a fixed center, which is not necessarily centroidal. Therefore, we need to adapt and give special treatment for $\sigma$ as an element of the optimal solution. Also, in the swapping iteration above, we never swap out $\sigma$, since the desired output must always contain $\sigma$, which also alters the swap pairs mapping as in Kanungo et al. (2002).
We adapt the mapping scheme of Kanungo et al. (2002), described in the Appendix with one modification: we set up a fixed capturing pair between $\sigma\in S$ and $\sigma\in O$. It makes sense since $\sigma=o_{\sigma}=s_{\sigma}$, as the closest center in $S$ to $\sigma$ is itself. With it, we always generate a swap partition of $S_{j}$ and $O_{j}$ that both contain $\sigma$. The reason we must have this swap partition (even though we never actually perform this swap on $\sigma$) is that we will sum up the cost occurred by every center in $S$ and $O$ by pair-by-pair from which the total costs $\Delta(S)$ and $\Delta(O)$ are formed (in Lemma 5). For this partition, there are two cases:
  • $|S_{j}|=|O_{j}|=1$, which means the partitions contain only $\sigma$, and hence when they are swapped, nothing changes from $S$ to $S^{\prime}$ and hence do not violate the constraint we set above about not swapping out $\sigma$.
  • $|S_{j}|=|O_{j}|>1$, which means the partitions contain some $o\in O$ that $s_{o}=\sigma$. By the design of the swap pairs mapping, all $s\in S_{j}\setminus\sigma$ will be chosen to be swapped out instead, and $\sigma\in S_{j}$ will never be swapped out.
For each swap pair of $s$ and $o$, we analyze the change in the distortion. For the insertion of $o$, all points in $N_{O}(o)$ are assigned to $o$, leads to the change of distortion to be: $\sum_{q\in N_{O}(o)}\big{(}\Delta(q,o)-\Delta(q,s_{q})\big{)}$.
For all points in $N_{S}(s)$ that lost $s$, they are assigned to new centers, which changes the distortion to be:
$\sum_{q\in N_{S}(s)\setminus N_{O}(o)}$ $\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)}\leq\sum_{q\in N_{S}(s)}\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)},$
and by the property $3$ of swap pairs, $s_{o_{q}}\neq s$, since $o_{q}\neq o$; the inequality is because for all $q$, $\Delta(q,s_{o_{q}})-\Delta(q,s)\geq 0$.
Lemma 5.
(Full proof in Lemma 15) Let $S$ be $1$-stable set and $O$ be the optimal set of $k$ centers, we have $\Delta(O)-3\Delta(S)+2R\geq 0$, where $R=\sum_{q\in P}\Delta(q,s_{o_{q}})$.
Lemma 6.
(Full proof in Lemma 17) With $R$ and $\alpha$ defined as above: $R\leq 2\Delta(O)+(1+2/\alpha)\Delta(S)$.
Proof of Theorem 4. By the results of Lemma 5 and 6:
00 0 $\leq\Delta(O)-3\Delta(S)+2R$ (12)
$\leq\Delta(O)-3\Delta(S)+4\Delta(O)+2\Delta(S)+\frac{4}{\alpha}\Delta(S)$ (13)
$=5\Delta(O)-(1-4/\alpha)\Delta(S).$ (14)
Using the same argument with Kanungo et al. (2002), we have $5/(1-4/\alpha)\geq\alpha^{2}$ or $(\alpha-5)(\alpha+1)\geq 0$ or $\alpha\leq 5$ which means $\frac{\Delta(S)}{\Delta(O)}=\alpha^{2}\leq 25$, and the Theorem follows.
In Section D.2, we describe how to create a $\gamma$-approximation candidate center set for any constant $\gamma$. Using it to create $C$ as a $\frac{25}{\gamma}$-approximation candidate center set and perform the above algorithm on $C$, the final output $S$ will be a $(25+\gamma)$-approximation.
Corollary 2.
Running peCluster with NonPrivateApproxFC be the above $k$-means algorithm, with probability at least $1-\beta$, $S^{(i)}_{\epsilon}$ is a $(w,t)$-approximation of $OPT_{i}$–the optimal $k$-means with a center fixed at position $x_{i}$, in which:
$w$ $=(25+\gamma)(1+\alpha)$
$t$ $=(25+\gamma)\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.polylog(n/\beta)/\epsilon\right).$
Figure 1: A detailed visualization of our dataset (Albemarle County, Virginia) and analysis includes (a) a Scatter plot of the full dataset with sampled points for contrastive analysis. (b) Comparison of k-median clustering with fixed and non-fixed centroids, both private and non-private. (c) Bar graph showing contrastive explanation differences for differential private and non-private k-median with a fixed centroid. (d) Heatmap indicating k-median error costs based on fixed centroid choices.
5 Experiments
In this part of our study, we focus on understanding the impact of the privacy budget $\epsilon$ on the trade-off between privacy and accuracy. Additionally, we explore how $\epsilon$ influences the quality of differentially private explanations.
To assess the impact of $\epsilon$ on privacy and accuracy, we will employ a set of metrics: 1) PO (Private Optimal) to evaluate the clustering cost using the differentially private algorithm, 2) PC (Private Contrastive) to determine the cost in fixed centroid differentially private clustering, 3) RO (Regular Optimal) to measure clustering cost in regular, non-private data, and 4) RC (Regular Contrastive) to analyze the cost of fixed centroid clustering in non-private data.
Measurements To measure the quality of explanations, we define two key metrics: 1) APE (Average Private Explanation), which calculates the difference between PC and PO to represent the decrease in utility for clustering as an explanatory output for agents, and 2) AE (Average Explanation), which is the difference between RO and RC, serving as a baseline to assess the extent to which differential privacy impacts the explanations. These metrics will enable us to effectively quantify the explanatory power of our approach.
We study the trade-offs between privacy and utility for the above metrics. The $\epsilon$ value balances privacy and utility: lower $\epsilon$ values prioritize privacy but may compromise utility.
Datasets Our study focuses on activity-based population datasets from Charlottesville City and Albemarle County, Virginia, previously used for mobile vaccine clinic deployment Mehrab et al. (2022). The Charlottesville dataset includes 33,000 individuals, and Albemarle County has about 74,000. Due to our algorithm’s intensive computational demands, we sampled 100 points for contrastive explanations.
Computational intensity of our algorithm stems primarily from the k-median algorithm we employ, which formulates the problem using linear programming. When the input size is large, processing times for this specific k-median algorithm can increase significantly. However, the implementation we use is one of the more straightforward approaches to executing k-median with a fixed centroid, and faster algorithms could potentially be applied to alleviate these concerns. For this study, we have opted not to focus on optimizing the computation time but rather to sample the data points to expedite our results. Additionally, we chose to sample 100 points because this number strikes a balance between computational efficiency and the clarity of the results for the reader. Importantly, the number of sampled points does not substantially impact the outcome of our analysis, allowing us to maintain the integrity of our findings while ensuring readability and accessibility for our audience.
Furthermore, our method is data agnostic and does not rely on the s-sparse feature for sample points. It is sufficiently adaptable to handle sample points from any distribution. This flexibility ensures that our approach is robust and versatile, capable of effectively managing diverse datasets regardless of their sparsity or the specific characteristics of their distribution.
Dimensiality It is important to note that our experiments were conducted on a 2D dataset. This choice was driven by the fact that the initial step of our method involves reducing dimensionality, as outlined in DimReduction (Algorithm 1). This process entails scaling down and normalizing the data to produce a lower-dimensional dataset. By implementing this step at the outset, we effectively avoid the challenges associated with managing high-dimensional datasets, thus simplifying the data preparation phase.
Experimental results. In this section, we focus solely on the Albemarle County dataset, with additional analyses and datasets presented in the Appendix. Figure 1(a) displays the dataset, highlighting the original data points with blue dots and the sampled points used for contrastive explanations with orange ’X’ markers. In Figure 1(b), we observe significant trends in the k-median clustering results for both private and non-private settings. Both the Private Optimal (PO) and Private Contrastive (PC) exhibit a downward trend in the private setting, indicating that as the $\epsilon$ budget increases, the error cost decreases, which underscores the trade-off between privacy and accuracy. The trends for non-private settings are depicted as straight lines, reflecting their independence from the privacy budget and are included for comparative purposes.
Our analysis shows a consistent gap between the Private Optimal (PO) and Private Contrastive (PC) across different levels of $\epsilon$, marking a stable contrastive error that does not fluctuate with changes in the privacy budget. This stability is pivotal as it supports our goal of providing dependable explanations to agents, regardless of the $\epsilon$ value, ensuring that our approach adeptly balances privacy considerations with the demand for meaningful explainability.
The distinction between PO and PC arises because PO is derived from implementing the optimal k-median/k-means algorithm, while PC results from enforcing a centroid’s placement at a specific agent’s location, inherently producing a less optimal outcome due to this constraint. Initially, agents are presented with the optimal outcome (PO). When an agent inquires why a centroid is not positioned at their specific location, we offer a contrastive explanation (PC), which is naturally less optimal than the original result as it evaluates the utility change from hypothetically placing a centroid at that specific location. In this scenario, a higher score signifies lower utility, emphasizing the balance between providing personalized explanations and adhering to privacy constraints.
Figure 1(c) displays the average private explanation (APE) and average explanation (AE) with variance across different explanations. The average values of APE across multiple runs are relatively stable, irrespective of the privacy budget, reinforcing the robustness of our methodology against the constraints imposed by differential privacy.
Lastly, the heatmap in Figure 1(d) provides insights into how the relative spatial positioning affects the contrastive explanation. It vividly shows that the greater the distance a fixed centroid is from the distribution center, the larger the contrastive error, further illustrating the intricate dynamics of our clustering approach.
6 Discussion
Our work explores the design of private explanations for clustering, particularly focusing on the k-median and k-means objectives for Euclidean datasets. We formalize this as the PrivEC problem, which gives a contrastive explanation to each agent equal to the increase in utility when a cluster centroid is strategically positioned near them. The accuracy bounds of our algorithm match private clustering methods, even as they offer explanations for every user, within a defined privacy budget.
The related work in this domain has seen the development of algorithms for contrastive explanations, but our contribution stands out by integrating differential privacy guarantees. Our experiments demonstrate the resilience of our approach. Despite the added layer of providing differential private explanations on top of differential private clustering, the quality of our explanations remains uncompromised.
Appendix A Related work: additional details
Our work considers differential privacy for explainable AI in general (XAI) and Multi-agent explanations (XMASE) in particular, focusing on post-hoc contrastive explanations for clustering.
Extensive experiments presented in Saifullah et al. [2022] demonstrate non-negligible changes in explanations of black-box ML models through the introduction of privacy. The findings in Patel et al. [2022] corroborate these observations regarding explanations for black-box feature-based models. These explanations involve creating local approximations of the model’s behavior around specific points of interest, potentially utilizing sensitive data. In order to safeguard the privacy of the data used during the local approximation process of an eXplainable Artificial Intelligence (XAI) module, the researchers have devised an innovative adaptive differentially private algorithm. This algorithm is designed to determine the minimum privacy budget required to generate accurate explanations effectively. The study undertakes a comprehensive evaluation, employing both empirical and analytical methods, to assess how the introduction of randomness inherent in differential privacy algorithms impacts the faithfulness of the model explanations.
Nguyen et al. [2023] considers feature-based explanations (e.g., SHAP) that can expose the top important features that a black-box model focuses on. To prevent such expose they introduced a new concept of achieving local differential privacy (LDP) in the explanations, and from that, they established a defense, called XRAND, against such attacks. They showed that their mechanism restricts the information that the adversary can learn about the top important features while maintaining the faithfulness of the explanations.
The analysis presented in Goethals et al. [2022] considers security concerning contrastive explanations. The authors introduced the concept of the ”explanation linkage attack”, a potential vulnerability that arises when employing instance-based strategies to derive contrastive explanations. To address this concern, they put forth the notion of k-anonymous contrastive explanations. Furthermore, the study highlights the intricate balance between transparency, fairness, and privacy when incorporating k-anonymous explanations. As the degree of privacy constraints is heightened, a discernible trade-off comes into play: the quality of explanations and, consequently, transparency are compromised.
Amongst the three types of eXplainable AI mentioned earlier, the maintenance of privacy during explanation generation incurs a certain cost. This cost remains even if an expense was previously borne during the creation of the original model. However, in our proposed methodology for generating contrastive explanations in clustering scenarios, once the cost of upholding differential privacy in the initial solution is paid, no additional expenses are requisite to ensure differential privacy during the explanation generation phase.
Closer to our application is the study that investigates the privacy aspects concerning contrastive explanations in the context of team formation Georgara et al. [2022]. In this study, the authors present a comprehensive framework that integrates team formation solutions with their corresponding explanations, while also addressing potential privacy concerns associated with these explanations. To accomplish this, the authors introduce a privacy breach detector (PBD) that is designed to evaluate whether the provision of an explanation might lead to privacy breaches. The PBD consists of two main components: (a) A belief updater (BU), calculates the posterior beliefs that a user is likely to form after receiving the explanation. (b) A privacy checker (PC), examines whether the user’s expected posterior beliefs surpass a specified belief threshold, indicating a potential privacy breach. However, the research is still in its preliminary stages and needs a detailed evaluation of the privacy breach detector.
Our contribution includes the development of comprehensive algorithms for generating contrastive explanations with differential privacy guarantees. We have successfully demonstrated the effectiveness of these algorithms by providing rigorous proof for their privacy guarantees and conducting extensive experiments that showcased their accuracy and utility. In particular, we have shown the validity of our private explanations for clustering based on the $k$-median and $k$-means objectives for Euclidean datasets. Moreover, our algorithms have been proven to have the same accuracy bounds as the best private clustering methods, even though they provide explanations for all users, within a bounded privacy budget. Notably, our experiments in the dedicated experiments section reveal that the epsilon budget has minimal impact on the explainability of our results, further highlighting the robustness of our approach.
There has been a lot of work on private clustering and facility location, starting with  Gupta et al. [2010], which was followed by a lot of work on other clustering problems in different privacy models, e.g., Huang and Liu [2018], Stemmer [2020], Stemmer and Kaplan [2018], Nissim and Stemmer [2018], Feldman et al. [2017]. Gupta et al. [2010] demonstrated that the additive error bound for points in a metric space involves an $O(\Delta k^{2}\log{n}/\epsilon)$ ) term, where $\Delta$ is the space’s diameter. Consequently, all subsequent work, including ours, assumes points are restricted to a unit ball.
Appendix B Additional proofs for PECluster
Theorem 5.
(Full version of Theorem 1) The solution $c$ and $\text{cost}{(S_{\epsilon})}$ released by Algorithm 3 are $\epsilon$-DP. For all clients $i$, the value $\text{cost}(S_{\epsilon})-\text{cost}{S^{(i)}_{\epsilon}}$ released by Algorithm 4 is $\epsilon$-$x_{i}$-exclusive DP.
Proof.
It follows that $\text{cost}(S_{\epsilon})$ is the direct results (using no private information) of $Y$, which is $\epsilon/2$-differentially private coreset. By the post-processing property, $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon$-DP).
$c$ is the output of DimReverse that is $\epsilon/2$-DP w.r.t. the input $(X_{1},X_{2},\ldots,X_{k})$. By composition, $c$ is $\epsilon$-DP since the input $(X_{1},X_{2},\ldots,X_{k})$ is calculated partially from $Y$, which is $\epsilon/2$-DP.
For each explanations $S^{(i)}_{\epsilon}$, let $D,D^{\prime}:D\setminus\{x_{i}\}\sim D^{\prime}\setminus\{x_{i}\}$, let $S^{(i)}_{\epsilon}(Y^{D},x_{i})$ be the value of $S^{(i)}_{\epsilon}$ with input dataset $D$ (and $Y^{D}$ as the private coreset of $D$ respectively) and a fixed center $x_{i}$, let $T=\{Y:S^{(i)}_{\epsilon}(Y,x_{i})\in S\}$, we have $D\sim D^{\prime}$ and therefore:
$\Pr[S^{(i)}_{\epsilon}(D,x_{i})\in S]$ $=\Pr[Y^{D}\in T]$ (15)
$\leq e^{\epsilon/2}\Pr[Y^{D^{\prime}}\in T]$ (16)
$=e^{\epsilon/2}\Pr[S^{(i)}_{\epsilon}(D^{\prime},x_{i})\in T],$ (17)
which implies that $\text{cost}(S^{(i)}_{\epsilon})-\text{cost}(S_{\epsilon})$ is $\epsilon$-$x_{i}$-exclusive DP, since $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon/2$-$x_{i}$-exclusive DP). ∎
Lemma 7.
(Full version of Lemma 1) With probability at least $1-\beta$, $(c_{1},\ldots,c_{k})$ is a $(w,t)$-approximation of $OPT$, where:
$w$ $=w^{\prime\prime}(1+\alpha),$ (18)
$t$ $=\mathcal{O}_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right)$ (19)
Proof.
This result derives from Theorem 48 of Ghazi et al. [2020] (From Eq(20) to Eq(24) of Ghazi et al. [2020]). ∎
Lemma 8.
(Full version of Lemma 2) Fix an $i$. If $OPT_{i}\geq w^{\prime\prime}(1+\alpha)OPT+t^{(i)}$, then with probability at least $1-2\beta$, cost($S_{\epsilon}$) and cost($S^{(i)}_{\epsilon}$) released by Algorithm  3 and  4 satisfies that $cost(S^{(i)}_{\epsilon})>cost(S_{\epsilon}$)
Proof.
By the result of Lemma 9, with probability $1-2\beta$ we have:
$cost(S^{(i)}_{\epsilon})$ $\geq OPT_{i}$ (20)
$\geq w^{\prime\prime}(1+\alpha)OPT+t^{(i)}$ (21)
$\geq w^{\prime\prime}(1+\alpha)\frac{cost(S_{\epsilon})-\Omega_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right)}{w^{\prime\prime}(1+\alpha)}+t^{(i)}$ (22)
$=cost(S_{\epsilon})+t^{(i)}-\Omega_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right).$ (23)
Lemma 9.
(Johnson-Lindenstrauss (JL) Lemma Johnson and Lindenstrauss [1984], Dasgupta and Gupta [2003]) Let $v$ be any $d$-dimensional vector. Let $\mathcal{S}$ denote a random $d^{\prime}$-dimensional subspace of $\mathbb{R}^{d}$ and let $\Pi_{S}$ denote the projection from $\mathbb{R}^{d}$ to $\mathcal{S}$. Then, for any $\zeta\in(0,1)$ we have
$\Pr\left[\|v\|_{2}\approx_{1+\zeta}\sqrt{d/d^{\prime}}\|\Pi_{\mathcal{S}}(v)\|_{2}\right]\geq 1-2\exp\left(-\frac{d^{\prime}\zeta^{2}}{100}\right)$ (24)
Lemma 10.
(Dimensionality Reduction for $(k,p)$-Cluster Makarychev et al. [2019]) For every $\beta>0,\tilde{\alpha}<1,p\geq 1,k\in\mathbb{N}$ , there exists $d^{\prime}=O_{\tilde{\alpha}}(p^{4}\log(k/\beta))$. Let $\mathcal{S}$ be a random d-dimensional subspace of $\mathbb{R}^{d}$ and $\Pi_{\mathcal{S}}$ denote the projection from $\mathbb{R}^{d}$ to $\mathcal{S}$. With probability $1-\beta$, the following holds for every partition $\mathcal{X}=(X_{1},\ldots,X_{k})$ of $X$:
$cost^{p}(\mathcal{X})\approx_{1+\tilde{\alpha}}(d/d^{\prime})^{p/2}cost^{p}(\Pi_{\mathcal{S}}(\mathcal{X})),$ (25)
where $\Pi_{\mathcal{S}}(\mathcal{X})$ denotes the partition $(\Pi_{\mathcal{S}}(X_{1}),\ldots,\Pi_{\mathcal{S}}(X_{k}))$.
Theorem 6.
(Full version of Theorem 2) Fix an $i$. With probability at least $1-\beta$, cost($S^{(i)}_{\epsilon}$) released by Algorithm 4 is a $(w,t)$-approximation of the optimal cost $OPT_{i}$, in which:
$w$ $=w^{\prime}(1+\alpha)$
$t$ $=w^{\prime}\mathcal{O}_{k,p}\left(\frac{2^{O_{p,\alpha}(d)}k^{2}\log^{2}{n}}{\epsilon}.\text{polylog}(\frac{n}{\beta}\right),$
and $OPT_{i}$ is the optimal $(k,p)$-clustering cost with a center is fixed at position $x_{i}$.
Proof.
Let $\tilde{\mathbf{X}}=(\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{n})$ and $X=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{n})$, setting $\alpha^{\prime}=0.1\alpha$, applying Lemma 10, we have:
${OPT}_{i}^{\tilde{d}}\leq\left(\frac{d^{\prime}}{d}\right)^{p/2}(1+0.1\alpha)OPT_{i}^{d}.$ (26)
By standard concentration, it can be proved that $\|x^{\prime}_{i}\|\leq 1/\Lambda$ with probability $0.1\beta/n$ as follows:
Using Lemma 9, we have:
$\Pr\left[\|x\|>\frac{1}{1+\zeta}\sqrt{d/d^{\prime}}\|x^{\prime}\|\right]$ $\geq 1-2\exp(-d^{\prime}\zeta^{2}/100).$ (27)
Since $x$ is in the unit ball, $\|x\|<1$, which leads to:
$\Pr\left[\|x^{\prime}\|<(1+\zeta)\sqrt{d^{\prime}/d}\right]$ $\geq 1-2\exp(-d^{\prime}\zeta^{2}/100).$ (28)
Setting $\zeta=\sqrt{\frac{log(n/\beta)}{0.01}}-1$, $\Lambda=\frac{1}{1+\zeta}\sqrt{d/d^{\prime}}=\sqrt{\frac{0.01}{\log{(n/\beta)}}.\frac{d}{d^{\prime}}}$, we have:
$\Pr[\|x^{\prime}\|<1/\Lambda]$ $\geq 1-2\exp(-\frac{d^{\prime}\zeta^{2}}{100})$ (29)
$>1-2\exp(-d^{\prime}\log(n/\beta))$ (30)
$>1-2\beta/n,$ (31)
By union bound on all $i$, then with probability at least $1-2\beta$, $x^{\prime}_{i}=\Lambda\tilde{x}_{i}$ for all $i$. Since $Y$ is the output of PrivateCoreset with input $X$, then by Theorem 38 of Ghazi et al. [2020], $Y$ is a $(0.1\alpha,t)$-coreset of $X^{\prime}$ (with probability at least $1-\beta$), with $t^{\prime}$ as:
$t^{\prime}=O_{p,\alpha}\left(\frac{2^{O_{p,\alpha}(d^{\prime})}k^{2}\log^{2}{n}}{\epsilon}\log\left(\frac{n}{\beta}\right)+1\right)$ (32)
Let $(y_{1},y_{2},\ldots,y_{k})$ be the solution of NonPrivateApproxFC in peCluster for a fixed $i$, $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ be the optimal solution of the clustering with fixed center at $x^{\prime}_{i}$ on $X$, $OPT_{Y}$ be the optimal cost of the clustering with fixed center at $x^{\prime}_{i}$ on $Y$. By the $w^{\prime}$-approximation property of NonPrivateApproximationFC, we have:
$cost_{Y}(y_{1},y_{2},\ldots,y_{k})$ $\leq w^{\prime}OPT_{Y}$ (33)
$\leq w^{\prime}cost_{Y}(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ (34)
$\leq w^{\prime}(1+0.1\alpha)cost_{\mathbf{X}^{\prime}_{1..n}}(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})+w^{\prime}t^{\prime}$ (35)
$=w^{\prime}(1+0.1\alpha)OPT^{d^{\prime}}_{i}+w^{\prime}t^{\prime}.$ (36)
Composing with Lemma 10, we have:
$cost_{Y}(y_{1},y_{2},\ldots,y_{k})$ $\leq w^{\prime}(1+0.1\alpha)OPT^{d^{\prime}}_{i}+w^{\prime}t^{\prime}$ (37)
$\leq\Lambda^{p}w^{\prime}(1+0.1\alpha)OPT^{\tilde{d}}_{i}+w^{\prime}t^{\prime}$ (38)
$\leq\Lambda^{p}w^{\prime}(1+0.1\alpha)(1+0.1\alpha)\left(\frac{d^{\prime}}{d}\right)^{p/2}OPT^{d}_{i}+w^{\prime}t^{\prime}$ (39)
$\leq w^{\prime}(1+\alpha)OPT^{d}_{i}\left(\frac{0.01}{\log(n/\beta)}\right)^{p/2}+w^{\prime}t^{\prime},\mbox{ since $\Lambda^{2}d^{\prime}/d=\Theta(1/\log(n/\beta))$}$ (40)
Finally, since
$cost(S^{(i)}_{\epsilon})$ $=cost_{Y}(y_{1},y_{2},\ldots,y_{k})\left(\frac{\log(n/\beta)}{0.01}\right)^{p/2}$ (41)
$\leq w^{\prime}(1+\alpha)OPT^{d}_{i}+\Theta(w^{\prime}t^{\prime}(\log(n/\beta)^{p/2})$ (42)
$\overset{(a)}{\leq}w^{\prime}(1+\alpha)OPT^{d}_{i}+w^{\prime}\Theta\left(\frac{2^{O_{p,\alpha}(d^{\prime})}k^{2}\log^{2}{n}}{\epsilon}\log\left(\frac{n}{\beta}\right)(\log(n/\beta))^{p/2}\right)$ (43)
$\overset{(b)}{\leq}w^{\prime}(1+\alpha)OPT^{d}_{i}+w^{\prime}\Theta\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}k^{2}\log^{2}{n}}{\epsilon}\log\left(\frac{n}{\beta}\right)(\log(n/\beta))^{p/2}\right)$ (44)
$\overset{(c)}{\leq}w^{\prime}(1+\alpha)OPT^{d}_{i}+w^{\prime}\Theta\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}(k/\beta)^{2}\log^{2}(n/\beta)}{\epsilon}\log\left(\frac{n}{\beta}\right)(\log(n/\beta))^{p/2}\right)$ (45)
$=w^{\prime}(1+\alpha)OPT^{d}_{i}+w^{\prime}\Theta\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right),$ (46)
where in $(a)$ we substitute the value of $t^{\prime}$, and $(b)$ is because $d^{\prime}=O_{\alpha}(p^{4}\log(n/\beta))$, and $(c)$ is because $\beta<1$, and the Lemma follows. ∎
Set $t^{(i)}=\Omega_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right)$ and the Lemma follows. ∎
Appendix C Additional proofs for $k$-median algorithm
Definition 9.
The solution of the $k$-median problem (with demands and a center fixed at a location $C$) can be formulated as finding the optimal solution of the following Integer program (IP):
minimize $\sum_{i,j\in N}d_{j}c_{ij}x_{ij}$ (48)
subject to $\text{subject to }\sum_{i\in N}x_{ij}$ $=1\text{ for each $j\in N$}$ (49)
$x_{ij}$ $\leq y_{i}\text{ for each $i,j\in N$}$ (50)
$\sum_{j\in N}y_{i}$ $=k$ (51)
$x_{ij}$$\in\{0,1\}\text{ for each $i,j\in N$}$ (52)
$y_{i}$$\in\{0,1\}\text{ for each $i\in N$}$ (53)
$y_{C}$ $=1\text{ for a fixed $C\in N$}$ (54)
$x_{CC}$ $=1\text{ for a fixed $C\in N$}.$ (55)
Lemma 11.
Locations $i,j\in N^{\prime}$ satisfy: $c_{ij}>4\max(\bar{C}_{i},\bar{C}_{j})$.
Proof.
The lemma follows the demands moving step (in the first step of the algorithm): for every $j$ to the right of $i$ (which means $\bar{C}_{j}\geq\bar{C}_{i}$) and within the distance of $\bar{C}_{j}$ (that also covers all points within distance $\bar{C}_{i}$), we move all demands of $j$ to $i$, hence $j$ will not appear in $N^{\prime}$. ∎
Lemma 12.
(Lemma 4 of Charikar et al. [1999]) For any feasible integer solution $(x^{\prime},y^{\prime})$ for the input with modified demands, there is a feasible integer solution for the original input with cost at most $4\bar{C}_{LP}$ plus the cost of $(x^{\prime},y^{\prime})$ with demands $d^{\prime}$.
Lemma 13.
(Theorem 6 of Charikar et al. [1999]) There is a $1/2$-restricted solution $(x^{\prime},y^{\prime})$ of cost at most $2\bar{C}_{LP}$.
Appendix D Additional proofs for $k$-means algorithm
Lemma 14.
(Lemma 2.1 of Kanungo et al. [2002]) Given a finite subset $S$ of points in $\mathbb{R}^{d}$, let $c$ be the centroid of $S$. Then for any $c^{\prime}\in\mathbb{R}^{d}$, $\Delta(S,c^{\prime})=\Delta(S,c)+|S|\Delta(c,c^{\prime})$.
Lemma 15.
(Full version of Lemma 5) Let $S$ be $1$-stable set and $O$ be the optimal set of $k$ centers, we have $\Delta(O)-3\Delta(S)+2R\geq 0$, where $R=\sum_{q\in P}\Delta(q,s_{o_{q}})$.
Proof.
Since $S$ is $1$-stable, we have for each swap pair:
$\sum_{q\in N_{O}(o)}$ $\big{(}\Delta(q,o)-\Delta(q,s_{q})\big{)}$ (56)
$+\sum_{q\in N_{S}(s)\setminus N_{O}(o)}$ $\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)}\geq 0.$ (57)
We will sum up the inequality above for all swap pairs. For the left term, the sum is overall $o\in O$:
$\sum_{o\in O}\sum_{q\in N_{O}(o)}\big{(}\Delta(q,o)-\Delta(q,s_{q})\big{)}$ (58)
$=$ $\sum_{q\in P}\big{(}\Delta(q,o)-\Delta(q,s_{q})\big{)},$ (59)
Since each $o\in O$ will appear exactly once, and $\cup_{o\in O}q\in N_{O}(o)$ will cover all points in $P$.
For the right term, the sum is over all $s$ that is being swapped out. We note that each $s$ can be swapped out at most twice, hence:
s $\sum_{s\text{ being swapped out}}\sum_{q\in N_{S}(s)}$ $\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)}$ (60)
$\leq 2\sum_{q\in P}$ $\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)}$ (61)
When we combine the two terms, we have:
$\sum_{q\in P}\big{(}\Delta(q,o)-\Delta(q,s_{q})\big{)}+2\sum_{q\in P}\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)}$ $\geq 0$ (62)
$\sum_{q\in P}\Delta(q,o_{q})-3\sum_{q\in P}\Delta(s,s_{q})+2\sum_{q\in P}\Delta(q,s_{o_{q}})$ $\geq 0$ (63)
$\Delta(O)-3\Delta(S)+2R$ $\geq 0,$ (64)
and the Lemma follows. ∎
Lemma 16.
(Proof in Lemma 2.2 & 2.3 of Kanungo et al. [2002]) Let $\alpha^{2}=\frac{\Delta(S)}{\Delta(O)}$, we have $\sum_{q\in P}dist(q,o_{q})dist(q,s_{q})\leq\frac{\Delta(S)}{\alpha}$
Lemma 17.
(Full version of Lemma 6) With $R$ and $\alpha$ defined as above: $R\leq 2\Delta(O)+(1+2/\alpha)\Delta(S)$.
Proof.
By Lemma 15, we have:
$R$ $=\sum_{q\in P}\Delta(q,s_{o_{q}})$ (65)
$=\sum_{o\in O}\sum_{q\in N_{O}(o)}\Delta(q,s_{o})$ (66)
$=\sum_{o\in O\setminus\sigma}\sum_{q\in N_{O}(o)}\Delta(q,s_{o})+\sum_{q\in N_{O}(\sigma)}\Delta(q,\sigma)$ (67)
$=\sum_{o\in O\setminus\sigma}\Delta(N_{O}(o),s_{o})+\Delta(N_{O}(\sigma),\sigma)$ (68)
=$\overset{(a)}{=}\sum_{o\in O\setminus\sigma}\big{(}\Delta(N_{O}(o),o)+|N_{O}(o)|\Delta(o,s_{o})\big{)}+\Delta(N_{O}(\sigma),\sigma)$ (69)
$=\sum_{o\in O\setminus\sigma}\sum_{q\in N_{O}(o)}\big{(}\Delta(q,o)+\Delta(o,s_{o})\big{)}+\sum_{q\in N_{O}(\sigma)}\Delta(q,\sigma)+\Delta(\sigma,o_{\sigma})$ (70)
$=\sum_{o\in O}\sum_{q\in N_{O}(o)}\big{(}\Delta(q,o)+\Delta(o,s_{o})\big{)}$ (71)
=$\overset{(b)}{=}\sum_{o\in O}\sum_{q\in N_{O}(o)}\big{(}\Delta(q,o)+\Delta(o,s_{q})\big{)}$ (72)
$\leq\sum_{q\in P}\big{(}\Delta(q,o_{q})+\Delta(o_{q},s_{q})\big{)}$ (73)
$\overset{(c)}{\leq}\Delta(O)+\sum_{q\in P}(dist(o_{q},q)+dist(q,s_{q}))^{2}$ (74)
$=2\Delta(O)+\Delta(S)+2\sum_{q\in P}dist(q,o_{q})dist(q,s_{q})$ (75)
$\overset{(d)}{\leq}2\Delta(O)+\Delta(S)+(2/\alpha)\Delta(S),$ (76)
where $(a)$ is because Lemma 14 applies for all $o\in O\setminus\sigma$, $(b)$ is because $\Delta(o,s_{o})\leq\Delta(o,s_{q})$, $(c)$ is because the triangle inequality applies for $\Delta(o_{q},s_{q})$ and $(d)$ is because of Lemma 16 and the Lemma follows. ∎
D.1 Swap pairs mapping
In this section, we describe the swap pairs mapping scheme for the $k$-means with a fixed center algorithm. We adapt the scheme of Matoušek [2000] to accommodate the fixed center. We discuss the modifications in Section 4.2. Here we discuss the complete mapping scheme.
At the last iteration of the algorithm, we always have a candidate set of centers $S$ that is $1$-stable, i.e., no single feasible swap can decrease its cost. We then analyze some hypothetical swapping schemes, in which we try to swap a center $s\in S$ with an optimal center $o\in O$. We utilize the fact that such single swaps do not decrease the cost to create some relationships between $\Delta(S)$ and $\Delta(O)$–the optimal cost. Particularly, these relationships are stated in Lemma 15 and Lemma 17.
Let $\sigma$ be the fixed center. We note that $\sigma\in S$ and $\sigma\in O$. Let $s_{o}$ be the closest center in $S$ for an optimal center $o\in O$, which means $o$ is captured by $s_{o}$. It follows that $s_{\sigma}=\sigma$. A center $s\in S$ may capture no optimal center (we call it lonely). We partition both $S$ and $O$ into $S_{1},\ldots,S_{r}$ and $O_{1},\ldots,O_{r}$ that $|S_{i}|=|O_{i}|$ for all $i$.
We construct each pair of partitions $S_{i},O_{i}$ as follows: let $s$ be a non-lonely center, $O_{i}=\{o\in O:s_{o}=s\}$, i.e., $O_{i}$ is the set of all optimal centers that are captured by $s$. Now, we compose $s$ with $|O_{i}|-1$ lonely centers (which are not partitioned into any group from $S$) to form $S_{i}$. It is clear that $|S_{i}|=|O_{i}|\geq 1$.
We then generate swap pairs for each pair of partitions $S_{i},O_{i}$ by the following cases:
  • $|S_{i}|=|O_{i}|=1$: let $S_{i}=\{s\},O_{i}=\{o\}$, generate a swap pair $\{s,o\}$.
  • $|S_{i}|=|O_{i}|=m>1$: let $S_{i}=\{s,s_{1..m-1}\}$ in which $s_{1..m-1}$ are $m-1$ lonely centers, let $O_{i}=\{o_{1..m}\}$, generate $m-1$ swap pairs $\{s_{j},o_{j}\}$ for $j=1..m-1$. Also, we generate a swap pair of $\{s_{1},o_{m}\}$. Please note that $s$ does not belong to any swap pair, each $o_{j}$ belongs to exactly one swap pair, and each $s_{j}$ belongs to at most two swap pairs.
We then guarantee the following $3$ properties of our swap pairs:
  1. 1.
    each $o\in O$ is swapped in exactly once
  2. 2.
    each $s\in S$ is swapped out at most twice
  3. 3.
    for each swap pair $\{s,o\}$, $s$ either captures only $o$, or $s$ is lonely (captures nothing).
D.2 $\gamma$-approximate candidate center set for fixed-center $k$-means.
We describe how to generate a $\gamma$-approximate candidate center set for $k$-means with fixed center $\sigma$ for a dataset $X\subset\mathbb{R}^{d}$. From the result of Matoušek [2000], we create a set $C^{\prime}$ which is a $\gamma$-approximation centroid set of $X$. We will prove that $C=C^{\prime}\cup\{\sigma\}$ forms a $\gamma$-approximate candidate center set for $k$-means with fixed center $\sigma$.
Definition 10.
Let $S\subset\mathbb{R}^{d}$ be a finite set with its centroid $c(S)$. A $\gamma$-tolerance ball of $S$ is the ball centered at $c(S)$ and has radius of $\frac{\gamma}{3}\rho(S)$.
Definition 11.
Let $X\subset\mathbb{R}^{d}$ be a finite set. A finite set $C^{\prime}\in\mathbb{R}^{d}$ is a $\gamma$-approximation centroid set of $X$ if $C^{\prime}$ intersects the $\gamma$-tolerance ball of each nonempty $S\subseteq X$.
Lemma 18.
(Theorem 4.4 of Matoušek [2000]) We can compute $C^{\prime}$–a $\gamma$-approximation centroid set of $X$ that has size of $O(n\gamma^{-d}\log(1/\gamma))$ in time $O(n\log{n}+n\gamma^{-d}\log(1/\gamma))$ ).
Theorem 7.
Let $C=C^{\prime}\cup\{\sigma\}$, in which $C^{\prime}$ is a $\gamma$-approximation centroid set computed as Lemma 18, then $C$ is a $\gamma$-approximate candidate center set for $k$-means with fixed center $\sigma$.
Proof.
Let $O=(O_{1},O_{2},\ldots,O_{k})$ be the optimal clustering in which $O_{1}$ is the cluster whose center is $\sigma$ (we denote it as $O_{\sigma}$). For any $S\subset\mathbb{R}^{d}$, we define $cost_{S}(c)=\sum_{x\in S}\|x-c\|^{2}$ and $cost(S)=cost_{S}(c(S))$ in which $c(S)$ is the centroid of $S$. By Definition 7, we will prove that there exists a set $c_{1},c_{2},\ldots,c_{k}\subset C$ and $c_{1}=\sigma$ such that $cost(c_{1},c_{2},\ldots,c_{k})\leq(1+\gamma)cost(O)$. We adapt the analysis of Matoušek [2000] for the special center $\sigma$–which is not a centroid as other centers in $k$-means.
First, we analyze the optimal cost. For any cluster except $O_{\sigma}$, its center is also the centroid $c(O_{i})$ of the cluster, while $O_{\sigma}$ must have center $\sigma$:
$cost(O)$ $=\sum_{x\in O_{\sigma}}\|x-\sigma\|^{2}+\sum_{i=2..k}\sum_{x\in O_{i}}\|x-c(O_{i})\|^{2}$ (77)
$=cost_{O\sigma}(\sigma)+\sum_{i=2..k}cost(O_{i})$ (78)
Now, we construct $\{c_{1},\ldots,c_{k}\}$ as follows: setting $c_{1}=\sigma$, for $i=2..k$, $c_{i}\in C^{\prime}$ is the candidate center that intersects the $\gamma$-tolerance ball of cluster $O_{i}$. For $O_{\sigma}$, $cost_{O_{\sigma}}(\sigma)=cost(O_{\sigma})$. For other clusters, $cost_{O_{i}}(c_{i})\leq(1+\gamma)cost(O_{i})$ as below:
$cost_{O_{i}}(c_{i})$ $=\sum_{x\in O_{i}}\|x-c_{i}\|^{2}$ (79)
$\leq\sum_{x\in O_{i}}(\|x-c(O_{i})\|+\|c(O_{i})-c_{i}\|)^{2}$ (80)
$=cost(O_{i})+2\|c_{i}-c(O_{i})\|\sum_{x\in O_{i}}\|x-c(O_{i})\|+|O_{i}|\|c_{i}-c(O_{i})\|^{2}$ (81)
$\leq cost(O_{i})+2\gamma/3\rho(O_{i})\sqrt{|O_{i}|}\sqrt{cost(O_{i})}+|O_{i}|(\gamma/3\rho(O_{i}))^{2}$ (82)
$\leq cost(O_{i})+(2/3)\gamma cost(O_{i})+(\gamma^{2}/9)cost(O_{i})$ (83)
$\leq(1+\gamma/3)^{2}cost(O_{i})$ (84)
$\leq(1+\gamma)cost(O_{i}).$ (85)
Let $(S_{1},S_{2},\ldots,S_{k})$ be the Voronoi partition with centers $(c_{1},c_{2},\ldots,c_{k})$, i.e., $S_{i}$ are points in the Voronoi region of $c_{i}$ in the Voronoi diagram created by $c_{1},\ldots,c_{k}$, we have:
$cost(c_{1},c_{2},\ldots,c_{k})$ $=cost_{S_{1}}(\sigma)+\sum_{i=2..k}cost(S_{i})$ (86)
$\overset{(a)}{\leq}cost_{S_{1}}(\sigma)+\sum_{i=2..k}cost_{S_{i}}(c_{i})$ (87)
$\overset{(b)}{\leq}cost_{O_{\sigma}}(\sigma)+\sum_{i=2..k}cost_{O_{i}}(c_{i})$ (88)
$\overset{(c)}{\leq}cost_{O_{\sigma}}(\sigma)+(1+\gamma)\sum_{i=2..k}cost(O_{i})$ (89)
$\leq(1+\gamma)cost(O),$ (90)
where $(a)$ is because $cost(S_{i})$ implies its minimal cost for any center, $(b)$ is because $S_{i}$s are picked by Voronoi partition which minimies the cost over $k$ partitions of seletecd $k$ centers, and $(c)$ is because $cost_{O_{i}}(c_{i})\leq(1+\gamma)cost(O_{i})$ as we proved above, and the Theorem follows. ∎
Figure 2: A detailed visualization of our dataset (Charlottesville County, Virginia) and analysis includes (a) a Scatter plot of the full dataset with sampled points for contrastive analysis. (b) Comparison of $k$-median clustering with fixed and non-fixed centroids, both private and non-private. (c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-median with a fixed centroid. (d) Heatmap indicating $k$-median error costs based on fixed centroid choices.
Figure 3: A detailed visualization of our synthetic dataset, measured by the same metrics as the other real-world datasets.
D.3 Additional details on experiments
D.3.1 Datasets and Experimental Setup
Beyond the Albemarle County dataset discussed in the main body of the paper, we also conducted experiments on other datasets, including the Charlottesville city dataset and synthetic activity-based populations. Here, we provide a detailed analysis of these experiments.
Charlottesville City Dataset:
This dataset is part of the synthetic U.S. population, as described in  Chen et al. [2021], Barrett et al. [2009]. It comprises about 33K individuals and approximately 5600 activity locations visited by these individuals. The locations represent places where individuals perform various activities.
Synthetic 2D Dataset:
The synthetic dataset was meticulously crafted to emulate the distribution and characteristics of real-world datasets, ensuring a balance between realism and controlled variability. Comprising 100 data points in a 2D space, this dataset was uniformly distributed, mirroring the range observed in the real datasets we analyzed.
The primary motivation behind this synthetic dataset is to provide a sandbox environment, free from the unpredictable noise and anomalies of real-world data. This controlled setting is pivotal in understanding the core effects of differential privacy mechanisms, isolating them from external confounding factors. The dataset serves as a foundational tool in our experiments, allowing us to draw comparisons and validate our methodologies before applying them to more complex, real-world scenarios.
D.3.2 Experimental Results
Impact of $\epsilon$ on Private Optimal (PO) and Private Contrastive (PC):
For the Charlottesville city dataset, we observed similar trends to those in the Albemarle county dataset. As the $\epsilon$ value increased, prioritizing accuracy, there was a slight compromise in privacy. However, consistent with our hypothesis, the influence of the epsilon budget on the explainability of our outcomes was negligible.
Effect of Privacy on Explainability:
The contrastive explanations for the Charlottesville city dataset also showcased the resilience of our approach. Even with differential privacy perturbations, the quality of explanations remained consistent, emphasizing the robustness of our method.
Visualization:
Figure  2 - (a) showcases the Charlottesville city dataset. Blue dots represent the original dataset, while orange ’X’ markers indicate the sampled points for contrastives. The distribution of these points provides a visual representation of the data’s diversity and the regions we focused on for our contrastive analysis.
D.3.3 Performance Evaluation:
For each $\epsilon$ value, we conducted 25 different runs on the Charlottesville city dataset. The average results were consistent with our findings from the Albemarle County dataset. It’s essential to note that these multiple invocations were solely for performance evaluation. In real-world applications, invoking private algorithms multiple times could degrade the privacy guarantee.
Consistency in Contrastive Explanations across Datasets:
Despite the distinct scales between the Charlottesville city and the Albemarle county datasets, we observed consistent patterns in the contrastive explanations. Specifically, as illustrated in Figure 2 - (b) and Figure  2 - (c), the contrastive explanations remained largely unaffected by variations in the $\epsilon$ budget. This consistency further reinforces our hypothesis that the epsilon budget has a negligible influence on the explainability of our outcomes, even when applied to datasets of different scales.
Impact of Fixed Centroids on Contrastive Error:
In Figure  2 - (d), we delve deeper into the effects of selecting different centroids as fixed centroids. It’s evident that the choice of a fixed centroid has a pronounced impact on the contrastive error. This underscores the importance of the centroid’s position concerning the data distribution. Some centroids, especially those closer to data-dense regions, tend to produce lower counterfactual errors, while others, particularly those near sparser regions or outliers, result in higher errors. This observation highlights the intricate relationship between centroid positioning and the resulting contrastive explanation’s accuracy.
D.3.4 Conclusion:
The extended experiments on the Charlottesville city dataset further validate our approach’s efficacy. The balance between privacy and utility, the robustness of contrastive explanations, and the negligible impact of $\epsilon$ on explainability were consistent across datasets. These findings underscore the potential of our method for diverse real-world applications.

### v2

@@
 Abstract
-This paper presents a novel approach in Explainable AI (XAI), integrating contrastive explanations with differential privacy in clustering methods.
-For several basic clustering problems, including $k$-median and $k$-means, we give efficient differential private contrastive explanations that achieve essentially the same explanations as those that non-private clustering explanations can obtain.
-We define contrastive explanations as the utility difference between the original clustering utility and utility from clustering with a specifically fixed centroid.
-In each contrastive scenario, we designate a specific data point as the fixed centroid position, enabling us to measure the impact of this constraint on clustering utility under differential privacy.
-Extensive experiments across various datasets show our method’s effectiveness in providing meaningful explanations without significantly compromising data privacy or clustering utility.
-This underscores our contribution to privacy-aware machine learning, demonstrating the feasibility of achieving a balance between privacy and utility in the explanation of clustering tasks.
+This paper presents a novel approach to Explainable AI (XAI) that combines contrastive explanations with differential privacy for clustering algorithms.
+Focusing on k-median and k-means problems, we calculate contrastive explanations as the utility difference between original clustering and clustering with a centroid fixed to a specific data point.
+This method provides personalized insights into centroid placement.
+Our key contribution is demonstrating that these differentially private explanations achieve essentially the same utility bounds as non-private explanations.
+Experiments across various datasets show that our approach offers meaningful, privacy-preserving, and individually relevant explanations without significantly compromising clustering utility.
+This work advances privacy-aware machine learning by balancing data protection, explanation quality, and personalization in clustering tasks.
 1 Introduction
 Different notions of clustering are fundamental primitives in several areas, including machine learning, data science, and operations research Reddy (2018).
-Clustering objectives (e.g., $k$-median and $k$-means), crucial in facility location planning, have been notably applied in pandemic response strategies, such as determining optimal sites for COVID-19 vaccination and testing centers  Bertsimas et al.
-(2022); Mehrab et al.
-(2022).
-The distance to a cluster center becomes a significant factor in these scenarios, particularly highlighted during the COVID-19 pandemic, where accessible vaccine locations played a pivotal role  Rader et al.
-(2022).
-This context often leads to individuals, especially those distant from a cluster center, seeking explanations for their specific assignment or location in the clustering plan.
-For example, consider a resident questioning why a COVID-19 testing center isn’t located next to their home.
-Such queries reflect the need for explainability in clustering decisions, where factors like population density, resource allocation, and public health considerations impact site selection.
-The resident’s inquiry underscores the importance of transparent methodologies in clustering that can provide understandable explanations for facility placements, especially in critical health scenarios.
+$k$-means and $k$-median clustering remain among the most important and widely used approaches, as demonstrated by recent advances in explainability, privacy, fairness, and contrastive learning Ghadiri et al.
+(2021); Moshkovitz et al.
+(2020); Zehtabi et al.
+(2024); Newling and Fleuret (2016).
+These problems often involve significant trade-offs between accessibility, resource allocation, and overall cost.
+For example, in emergency response planning, authorities must decide the optimal locations for ambulance stations to minimize response times across a city.
+Residents might question why an ambulance station isn’t located closer to their neighborhood, especially if they feel imbalances in resource distribution.
+Similarly, in retail, customers might wonder why a new store is not placed near their area, despite being part of a high-demand demographic.
+Explainability in these contexts provides transparency into how and why certain decisions are made, addressing questions like: ”Why was this location chosen instead of another?” This is particularly important in applications where the consequences of clustering decisions directly affect individuals or communities.
+Bobek et al.
+(2022); Ofek and Somech (2024)
 Such questions fall within the area of Explainable AI, which is a rapidly growing and vast area of research Finkelstein et al.
@@
 (2020); Miller (2019); Boggess et al.
+(2023); Schleibaum et al.
 (2024).
-We focus on post-hoc explanations, especially contrastive explanations, e.g., Miller (2019), which address “why P instead of Q?” questions.
-Such methods have been used in many applications, including multi-agent systems, reinforcement learning, and contrastive analysis, e.g., Boggess et al.
-(2023); Sreedharan et al.
+We focus on post-hoc explanations, especially contrastive explanations, e.g., Miller (2019); Pozanco et al.
+(2022), which address “why P instead of Q?” questions.
+For example, in warehouse optimization, contrastive explanations clarify why a specific location was chosen as a distribution center, considering constraints like storage capacity or demand Zehtabi et al.
+(2024).
+These methods are widely applied in multi-agent systems, reinforcement learning, and contrastive analysis Boggess et al.
+(2023, 2022); Sreedharan et al.
 (2021); Madumal et al.
@@
 (2020).
-In the context of the clustering problems mentioned above, a possible explanation for a resident is the increase in the cost of the clustering if a center is placed close to the resident, compared to the cost of an optimal solution, which is relatively easy to compute.
-This increased cost reflects a trade-off, where optimizing for one resident’s convenience could result in worse centroid positions for other residents, highlighting the complex decision-making process in clustering for fair resource distribution.
+In reinforcement learning, they explain actions by highlighting trade-offs, such as long-term rewards or risks van der Waa et al.
+(2018).
+Following the approach introduced in Zehtabi et al.
+(2024); Pozanco et al.
+(2022), we explain clustering decisions by comparing the costs in two scenarios: $cost(S)$, the cost of an optimal clustering solution on the whole dataset $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ and $cost(S^{(i)})$, the cost of a modified solution where we fix a center at a desired location requested by agent $x_{i}\in\mathbf{X}$.
+We explain the decision by showing $cost(S^{(i)})-cost(S)$, i.e., how much the overall clustering cost increases when we force a center to be in a specific location.
+A higher clustering cost indicates worse performance.
+This comparison reveals the trade-offs in clustering: optimizing for one specific location often leads to a higher overall cost, meaning a worse solution for everyone else.
+By examining the difference between the optimal clustering cost and the cost of the forced fixed clustering, we can understand why the algorithm chose certain locations for centers and not others.
+This approach helps people grasp the complex balancing act involved in clustering decisions, especially when trying to distribute resources fairly Zehtabi et al.
+(2024).
 Data privacy is a crucial concern across various fields, and Differential Privacy (DP) is one of the most widely used and rigorous models for privacy Dwork and Roth (2014).
-We focus on the setting where the set of data points $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ are private; for instance, in the vaccine center deployment problem, each $x_{i}$ corresponds to an individual who is seeking vaccines, and would like to be private.
-There has been a lot of work on the design of differentially private solutions to clustering problems like $k$-median and $k$-means under such a privacy model Ghazi et al.
+We focus on the setting where the set of data points $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ are private; for instance, in the ambulance center deployment problem, each $x_{i}$ represents an individual requiring emergency services and seeking to keep their information private.
+There has been a lot of work on the design of differentially private solutions to clustering problems such as $k$-median and $k$-means in such a privacy model Ghazi et al.
 (2020); Gupta et al.
@@
 (2017); Stemmer and Kaplan (2018).
-These methods guarantee privacy of the data points $\mathbf{X}$ and output a set of centers $c_{1},\ldots,c_{k}$ (which are public) such that its cost, denoted by $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})$ is within multiplicative and additive factors $w$ and $t$ of the optimal solution $OPT_{\mathbf{X}}$, i.e., $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})\leq w\cdot OPT_{\mathbf{X}}+t$.
 While there has been significant progress in various domains of differential privacy, the intersection of explainability and differential privacy still needs to be explored.
-In clustering problems, a potential private contrastive explanation given to agent $i$ is to provide the $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$, where $S_{\epsilon}$ is a set of centers generated by private algorithm (with privacy budget $\epsilon$), and $\text{cost}(S^{(i)}_{\epsilon})$ is the cost of the clustering solution while fixing one center at agent $i$’s location.
-However, giving such a private contrastive explanation to each agent $i$ naively using a $(w,t)$-approximate private clustering would lead to $(w,t\sqrt{n}/\epsilon)$-approximation, even if we use advanced composition techniques for privacy.
-The central question of our research: is it possible to offer each user a private contrastive explanation with a limited overall privacy budget?
+In clustering problems, building on the formalization of explanations for combinatorial problems we provide a private contrastive explanation to agent $x_{i}$ by computing $\text{cost}(S^{(i)}_{\epsilon})-\text{cost}(S_{\epsilon})$.
+Here, $S_{\epsilon}$ represents a private solution using the privacy budget $\epsilon$, while $S^{(i)}_{\epsilon}$ denotes a private solution, when we constrain one center to be at a location specified by agent $x_{i}$ (this can be any point of interest, not necessarily the agent’s own location).
+This difference quantifies how the clustering cost changes when accommodating agent $i$’s position, offering a privacy-preserving explanation of the clustering decision.
+However, giving such a private contrastive explanation to each agent $i$ naively using a private clustering algorithm would require a high privacy budget due to composition, which impact the accuracy, and lead to misleading or uninformative results.
+The central question of our research: is it possible to offer each user an informative private contrastive explanation with a limited overall privacy budget?
 Our contributions.
@@
 2.
-We present an $\epsilon$-DP mechanism, PrivateExplanations, which provides a contrastive explanation to each agent while ensuring the same utility bounds as private clustering in Euclidean spaces Ghazi et al.
-(2020).
+We present an $\epsilon$-DP mechanism, PrivateExplanations, which provides a contrastive explanation to each agent while ensuring the same utility bounds as private clustering in Euclidean spaces, offering personalized insights without compromising privacy or clustering quality.
 We use the private coreset technique of Ghazi et al.
 (2020), which is an intermediate private data structure that preserves similar clustering costs as the original data.
-The main technical challenge is to derive rigorous bounds on the approximation factors for all the contrastive explanations.
 3.
-We evaluate our methods empirically on realistic datasets used in the problems of vaccine center deployment in Virginia Mehrab et al.
-(2022).
-Our results show that the privacy-utility trade-offs are similar to those in private clustering, and the errors are quite low even for reasonable privacy budgets, demonstrating the effectiveness of our methods.
+We evaluate our methods on diverse datasets with varying distributions and feature dimensions.
+Our results demonstrate privacy-utility trade-offs comparable to private clustering, with low clustering errors even at reasonable privacy budgets, showcasing the effectiveness of our approach.
 Our research stands out by seamlessly integrating differential privacy into contrastive explanations, maintaining the quality of explanations even under privacy constraints.
 This work bridges the gap between privacy and explainability, marking a significant advancement in privacy-aware machine learning.
-Due to space limitations, we present many technical details and experimental evaluations in the Appendix.
+A key technical contribution of our work is the derivation of rigorous bounds on the approximation factors for all contrastive explanations, ensuring their reliability and effectiveness.
+Due to space limitations, we only present major technical details in the main paper.
+We maintain a full, updated version of this paper with complete proofs and extended experimental results at Nguyen et al.
+(2024).
 2 Related Work
@@
 We summarize some of the work directly related to our paper; additional discussion is presented in the Appendix, due to space limitations.
-Extensive experiments presented in Saifullah et al. (2022) demonstrate non-negligible changes in explanations of black-box ML models through the introduction of privacy.
+Extensive experiments presented in Saifullah et al.
+(2022) demonstrate non-negligible changes in explanations of black-box ML models through the introduction of privacy.
 Nguyen et al.
@@
 3 Preliminaries
-Let $\mathbf{X}\subset\mathbb{R}^{d}$ denote a dataset consisting of $d$-dimensional points (referred to sometimes as agents).
-We consider the notion of $(k,p)$-clustering, as defined by Ghazi et al.
-(2020).
+Let $\mathbf{X}\subset\mathbb{R}^{d}$ denote a dataset consisting of $d$-dimensional points (referred as agents). We consider the notion of $(k,p)$-clustering, as defined by Definition 1.
 Definition 1.
@@
 ($(w,t)$-approximation).
-Given $k\in\mathbb{N}$, and a multiset $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ of points in the unit ball, let $OPT^{p,k}_{\mathbf{X}}=\min_{c_{1},\ldots,c_{k}\in\mathbb{R}^{d}}\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})$ denote the cost of an optimal $(k,p)$-clustering.
+Given $k\in\mathbb{N}$, and a multiset $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ of points in the unit ball, let $OPT^{p,k}_{\mathbf{X}}=\min_{c_{1},\ldots,c_{k}\in\mathbb{R}^{d}}\\ \text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})$ denote the cost of an optimal $(k,p)$-clustering.
 We say $c_{1},\ldots,c_{k}$ is a $(w,t)$– approximation to a $(k,p)$-optimal clustering if $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})\leq w\cdot OPT^{p,k}_{\mathbf{X}}+t$.
+Let $OPT$ denote the cost of the optimal $(k,p)$-clustering, $OPT_{i}$ denote the cost of the optimal $(k,p)$-clustering, with a center fixed at position $z_{i}$ (the location chosen by agent $i$) and the remaining $k$-1 centers are chosen to optimize the objective.
+Let $w^{\prime},w^{\prime\prime}$ denote the maximum approximation (w.r.t.
+$OPT$ and $OPT_{i}$ respectively) of non-private clustering algorithms.
+These factors will be specified in Sections 4.1 and 4.2.
 A coreset (of some original set) is a set of points that, given any $k$ centers, the cost of clustering of the original set is “roughly” the same as that of the coreset Ghazi et al. (2020).
@@
 If $\delta=0$, we say $\mathcal{M}$ is $\epsilon$-differentially private.
-We operate under the assumption that the data points in $\mathbf{X}$ represent the private information of users or agents.
-We say $\mathbf{X},\mathbf{X}^{\prime}$ are neighboring (denoted by $\mathbf{X}\sim\mathbf{X}^{\prime}$) if they differ in one data point.
+We assume that the data points in $\mathbf{X}$ (i.e., users) are private, and say $\mathbf{X},\mathbf{X}^{\prime}$ are neighboring (denoted by $\mathbf{X}\sim\mathbf{X}^{\prime}$) if they differ in one data point.
 When a value is disclosed to an individual agent $i$, it is imperative to treat the remaining clients in $\mathbf{X}-\{i\}$ as private entities.
-To address this necessity, we adopt the following privacy model:
 Definition 5.
-A mechanism $\mathcal{M}$ is $\epsilon$-$i$-exclusive DP if, $\forall X,X^{\prime}:X\setminus\{i\}\sim X^{\prime}\setminus\{i\}$, and for all $S\subseteq Range(\mathcal{M})$:
+A mechanism $\mathcal{M}$ is $\epsilon$-$i$-exclusion DP if, $\forall X,X^{\prime}:i\in X,i\in X^{\prime},X\setminus\{i\}\sim X^{\prime}\setminus\{i\}$, and for all $S\subseteq Range(\mathcal{M})$:
 $\Pr[\mathcal{M}(X)\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(X^{\prime})\in S].$
-Note that if $\mathcal{M}$ is $\epsilon$-DP, it is $\epsilon$-$i$-exclusive DP for any $i$.
-Relation to Joint-DP model.
-Kearns et al.
-(2014) defines Joint-DP–a privacy model that “for each of agent $i$, knowledge of the $n-1$ agents does not reveal much about agent $i$”.
-It contrasts with our privacy model, where each explanation to agent $i$ uses direct and private knowledge of agent $i$ (that private knowledge will be returned to only agent $i$) and such knowledge does not reveal much about $n-1$ other agents.
-In clustering scenarios such as k-means or k-median, it is beneficial for an individual data point to be positioned near the cluster center, as this reduces its distance.
-This concept has practical applications in fields such as urban planning and public health.
-For example, clustering algorithms like $k$-means and $k$-median have been utilized to optimize the placement of mobile vaccine centers, as noted in  Mehrab et al.
+We extend this to say that $\mathcal{M}$ is $\epsilon$-$Y$-exclusion DP if the above holds $\forall X,X^{\prime}:Y\subset X,Y\subset X^{\prime},X\setminus Y\sim X^{\prime}\setminus Y$.
+We now define the PrivEC problem for providing private contrastive explanations, where each agent $x_{i}$ seeks an explanation for a center fixed at a location of their choosing denoted by $z_{i}$.
+Definition 6.
+Private and Explainable Clustering problem (PrivEC) Given an instance $\mathbf{X}\subset\mathbb{R}^{d}$, clustering parameters $k,p$, and a contrastive set of points $Z\subset\mathbb{R}^{d}$, the goal is to output:
+Private: An $\epsilon$-DP clustering solution $S_{\epsilon}$ (available to all)
+Explainable: For each agent $x_{i}\in X$, output $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$.
+$S^{(i)}_{\epsilon}$ is a private solution computed by the clustering algorithm with one centroid fixed at the position requested by agent $i$.
+We assume that $S^{(i)}_{\epsilon}$ is not revealed to any agent, but $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$ is released to agent $i$ as contrastive explanation, which is $\epsilon$-$i$-exclusion DP.
+Lemma 1.
+With probability at least $1-\beta$, cost($S_{\epsilon}$) (clustering cost) is a $(w,t)$-approximation of $OPT$, where [^1] We use the notation $\mathcal{O}_{p,\alpha}$ to explicitly ignore factors of $p,\alpha$:
+$w$ $=w^{\prime}(1+\alpha),$
+$t$ $=w^{\prime}\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.\text{polylog}(n/\beta)/\epsilon\right)$
+$\alpha$ is the approximation parameter for the utility of clustering and explanations. $\beta$ is the failure probability of the utility guarantees of clustering and explanations.
+Lemma 2.
+DP of fixed centroid yield additional cost.
+Fix an $i$.
+If $OPT_{i}\geq w^{\prime\prime}(1+\alpha)OPT+t^{(i)}$, then with probability at least $1-2\beta$, cost($S_{\epsilon}$) and cost($S^{(i)}_{\epsilon}$) computed by Algorithm peCluster satisfies that $cost(S^{(i)}_{\epsilon})>cost(S_{\epsilon}$).
+4 peCluster Mechanism
+We design PrivateExplanation (Algorithm 1) for providing contrastive explanations for each agent.
+Specifically, it takes as inputs: (1) $x_{i}$ which specifies the location of each agent $i$, and the contrastive location $z_{i}$ for which they want an explanation, (2) original and target dimensions (d, d’), number of clusters ($k$), privacy budget $\epsilon$, and $\zeta$ (explained later).
+The algorithm’s key components are:
+Algorithm 1 PrivateExplanation
+Input: $(x_{1},\ldots,x_{n}),(z_{1},\ldots,z_{n}),d,d^{\prime},k,\epsilon,\zeta$
+ Output: $(\epsilon,\delta)$-differentially private explanation for agents
+1:  $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$leftarrow $DimReduction((x_{1},\ldots,x_{n}),d,d^{\prime})$
+2:  $Z^{\prime}$leftarrow $DimReduction((z_{1},\ldots,z_{n}),d,d^{\prime})$
+3:  $Y\leftarrow\textsc{PrivateCoreset}^{\epsilon/2}(x^{\prime}_{1},\ldots,x^{\prime}_{n};\zeta)$
+4:  $(c^{\prime}_{1},\dots,c^{\prime}_{k}),cost(S^{\prime}_{\epsilon})\leftarrow NonPrivateApprox(Y,k)$
+5:  $cost(S_{\epsilon})=RevertDimValue*cost(S^{\prime}_{\epsilon})$
+6:  $c\leftarrow\textsc{DimReverse}^{\epsilon/2}((c^{\prime}_{1},\dots,c^{\prime}_{k}),(x^{\prime}_{1},\dots,x^{\prime}_{n}))$
+7:  for $z^{\prime}_{i}\in Z^{\prime}$ do
+8:     $cost(S^{\prime(i)}_{\epsilon})\leftarrow NonPrivateApproxFC(Y,k,z^{\prime}_{i})$
+9:     $cost(S^{(i)}_{\epsilon})=RevertDimValue*cost(S^{\prime(i)}_{\epsilon})$
+10:  end for
+11:  return ${cost(S^{(i)}_{\epsilon})-cost(S_{\epsilon})|i\in range(1,\ldots,|X|)}$
+  • Dimension Reduction: Using DimReduction from Ghazi et al.
+(2020), we transform input data (in dimension $d$) to a lower-dimensional space $d^{\prime}$.
+This reduction is crucial since our coreset algorithm is exponential in the dimension, but by reducing to logarithmic dimensions, it becomes polynomial-time.
+  • Private Coreset: We create a differentially private coreset $Y$ using PrivateCoreset from Ghazi et al. (2020), ensuring $\epsilon/2$ differential privacy; the coreset is defined in Definition 3.
+  • Clustering: The coreset is clustered using a non-private approximation algorithm (NonPrivateApprox).
+We can use a non-private clustering algorithm here since the coreset itself is already private, and by the Post-Processing property of differential privacy, the final result remains private.
+  • Cost Scaling: In line 5 of the algorithm, by multiplying by $\textsc{RevertDimValue}=(\log(n/\beta)/0.01)^{p/2}$ we scale the clustering cost ($cost(S^{\prime}_{\epsilon})$) (in the low-dimensional space) back to the original dimension (cost($S_{\epsilon}$)) as shown in Makarychev et al.
 (2019).
-However, once the cluster centers are calculated, residents might wonder why a vaccine clinic has not been positioned nearer to their homes.
-Our objective is to provide explanations for each data point that is not near a cluster center.
-Following the approach suggested by  Miller (2019), one way to provide an explanation is to re-run the clustering algorithm, this time setting the resident’s location as a fixed centroid.
-This allows us to explain the changes in clustering costs, where an increase could describe why certain placements are less optimal.
-The focus of this paper is twofold: residents not only need explanations but also wish to maintain their privacy, preferring not to disclose their precise locations/data publicly.
-Traditional private clustering methods struggle to offer these private explanations because they accumulate the privacy budget by continuously applying private clustering to answer numerous inquiries, ultimately resulting in explanations that are not private.
-Thus, we have developed a framework that delivers the necessary explanations to residents while protecting their confidentiality.
-In essence, our framework is data-agnostic and capable of offering contrastive explanations for any private data within clustering algorithms, effectively balancing the need for privacy with the demand for transparency in decision-making.
-Definition 6.
-Private and Explainable Clustering problem (PrivEC) Given an instance $\mathbf{X}\in\mathcal{X}^{n}$, clustering parameters $k,p$, and a subset of data points: $V_{s}\subset\mathbf{X}$, the goal is to output:
-Private: An $\epsilon$-DP clustering solution $S_{\epsilon}$ (available to all)
-Explainable: For each $i\in V_{s}$, output $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$, which is $\epsilon$-$i$-exclusive DP.
-$S^{(i)}_{\epsilon}$ is a private solution computed by the clustering algorithm while fixing one centroid to the position of agent $i$.
-We assume that $S^{(i)}_{\epsilon}$ is not revealed to any agent, but $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$ is released to agent $i$ as contrastive explanation.
-4 peCluster Mechanism
-The PrivateClustering algorithm Ghazi et al.
-(2020) includes dimension reduction, forming a private coreset, applying an approximation clustering algorithm on the coreset, and then reversing the dimension reduction, ultimately yielding an $\epsilon$-DP clustering solution.
-The PrivateExplanations algorithm receives the original private clustering cost from PrivateClustering and then leverages the PrivateClustering framework but introduces a fixed-centroid clustering algorithm on the coreset instead of regular approximation clustering algorithm, followed by a dimension reverse.
-Ultimately, after calculating the fixed centroid clustering algorithm and determining its cost in the original dimension, we subtract the original private clustering cost obtained from the PrivateClustering algorithm, outputting $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$.
-To elaborate further, Private Clustering first reduces the dataset’s dimensionality using DimReduction (Algorithm 1) to enable efficient differentially private coreset construction.
-It then uses a DP Exponential Mechanism to build a PrivateCoreset (Algorithm 7 of Ghazi et al.
-(2020)), which is clustered using any $w^{\prime\prime}$-approximate, not necessarily private clustering algorithm.
-The low-dimensional clustering cost is scaled to obtain the original high-dimensional cost, and the centroids are recovered in the original space.
-Multiplying the low-dimensional clustering cost with $(\log(n/\beta)/0.01)^{p/2}$ gives the clustering cost of the original high-dimensional instance Makarychev et al.
+This reversal is necessary because while we computed costs in reduced dimensions for efficiency, we need the final cost in the original dimensions for accuracy.
+  • Dimension Reverse: Centroids are mapped back to the original space using DimReverse, maintaining $\epsilon$-differential privacy.
+  • Contrastive Explanations: For each data point, we execute fixed-centroid clustering (NonPrivateApproxFC) on the coreset, constraining one centroid to a location chosen by the agent.
+This algorithm, detailed in Sections 4.1 and 4.2, is our key contribution as it modifies standard k-means and k-median algorithms to fix one centroid while maintaining their original utility bounds from literature, ensuring meaningful explanations.
+Without utility bounds, agents could challenge the validity of the explanation, arguing that the fixed centroid might degrade the clustering solution to an unacceptable extent.
+However, by guaranteeing the same utility bounds as the original algorithms, we ensure that the explanations are grounded in the vicinity of optimal clustering solutions, leaving no room for users to dispute the fairness or validity of the explanation.
+This alignment between explanation quality and clustering utility reinforces the trustworthiness of the algorithm and the insights it provides.
+After clustering, we apply RevertDimValue to transform the cost back to the original space (cost($S^{(i)}_{\epsilon}$)).
+By combining lines 5 and 9 of the algorithm, we derive the output: cost($S^{(i)}_{\epsilon}$) - cost($S_{\epsilon}$) for each agent.
+This value captures the loss of optimality when fixating one centroid, quantifying how much the clustering quality degrades due to this constraint, serving as a contrastive explanation.
+Theorem 1.
+DP of Explanation.
+The solution $(c_{1},\ldots,c_{k})$ and $cost(S_{\epsilon})$ computed by Algorithm peCluster are $\epsilon$-DP.
+For all clients $i$ and $S^{(i)}_{\epsilon}$ computed by Algorithm peCluster is $\epsilon$-$i$-exclusion DP.
+Privacy analysis, as demonstrated in Theorem 1, we establishes the privacy guarantees of peCluster.
+$Y$ coreset is $\epsilon/2$-differentially private as an output of $\epsilon/2$-DP algorithm.
+Consequently, $(c^{\prime}_{1},\ldots,c^{\prime}_{k})$ and $cost(S_{\epsilon})$ maintain $\epsilon/2$-DP status, under the Post-Processing property.
+Applying $\textsc{DimReverse}^{\epsilon/2}$ to find the centers in the original space, $c=\{c_{1},\ldots,c_{k}\}$ is $\epsilon$-DP by Composition theorem.
+For each $i$, cost($S^{(i)}_{\epsilon}$) is produced by Post-Processing of $Y$ with only $z^{\prime}_{i}$, hence $cost(S^{(i)}_{\epsilon}$) satisfies $\epsilon$-$i$-exclusion-DP.
+Running Time Analysis.
+Algorithm 1 has a total runtime of $O((k/\beta)^{O_{p,\alpha}(1)}\text{poly}(nd))$, which is polynomial in the input size.
+The key components contributing to this complexity include PrivateCoreset, DimReverse, and instances of $(k,p)$-clustering with and without fixed centers.
+PrivateCoreset runs in $O((k/\beta)^{O_{p,\alpha}(1)}\text{poly}(n))$ time, as it sets $d^{\prime}=O(p^{4}\log(k/\beta))$ to satisfy the Dimension-Reduction Lemma (Appendix Section B) and uses Lemma 42 from Ghazi et al.
 (2020).
-By Running DimReverse using FindCenter (Algorithm 10 of Ghazi et al.
-(2020)) we can find the centroids in the original space.
-The output is the private centroids, their cost, the private coreset, and the low-dimensional dataset.
-We employed DimReduction since PrivateCoreset is an exponential algorithm.
-This approach enables the transformation of dimensionality reduction into an exponential process, effectively converting the algorithm’s time complexity from exponential to polynomial with respect to the input size.
-The following summarizes Algo. 3 - PrivateClustering:
-Input: Agents’ location ($X$), original and reduced dimension (d, d’). Output: Private clustering: centers and cost.
-  1. 1.
-    Dimension reduction - Algorithm 1 DimReduction
-  2. 2.
-    Private coreset - Algorithm. PrivateCoreset
-  3. 3.
-    Clustering on the coreset - Algorithm NonPrivateApprox
-  4. 4.
-    Dimension reverse -  2 DimReverse
-Algorithm 1 DimReduction
-Input: $(x_{1},x_{2},\ldots,x_{n}),d,d^{\prime},\beta$
-Output: $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$ low-dimensional space dataset.
-1:  $\Lambda=\sqrt{\frac{0.01d}{\log(n/\beta)d^{\prime}}}$
-2:  for $i\in\{1,..,n\}$ do
-3:     $\tilde{x}_{i}\leftarrow\Pi_{\mathcal{S}}(x_{i})$
-4:     if $\|\tilde{x}_{i}\|\leq 1/\Lambda$ then
-5:        $x^{\prime}_{i}=\Lambda\tilde{x}_{i}$
-6:     else
-7:        $x^{\prime}_{i}=0$
-8:     end if
-9:  end for
-10:  return $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
-Algorithm 2 DimReverse
-Input: $(c^{\prime}_{1},\ldots,c^{\prime}_{k}),(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
-Output: $(c_{1},\ldots,c_{k})$ Private Centroids in high dimension
-1:  $\mathbf{X}_{1},\ldots,\mathbf{X}_{k}\leftarrow\text{ the partition induced by }$ $(c^{\prime}_{1},\ldots,c^{\prime}_{k})$ on $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
-2:  for $j\in\{1,\ldots,k\}$ do
-3:     $c_{j}\leftarrow\textsc{FindCenter}^{\epsilon/2}(\mathbf{X}_{j})$
-4:  end for
-5:  return $(c_{1},\ldots,c_{k})$
-Algorithm 3 PrivateClustering
-Input: $(x_{1},x_{2},\ldots,x_{n}),d,d^{\prime},k,\epsilon,\lambda,p,\alpha,\beta$
- Output: $(\epsilon,\delta)$-differentially private centers $c_{1,..,k}$ with cost $S_{\epsilon}$, $Y$ Private Coreset and $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$ low-dimensional space dataset.
-1:  $\zeta=0.01\left(\frac{\alpha}{10\lambda_{p,\alpha/2}}\right)^{p/2}$
-2:  $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$leftarrow $DimReduction((x_{1},\ldots,x_{n}),d,d^{\prime},\beta)$
-3:  $Y\leftarrow\textsc{PrivateCoreset}^{\epsilon/2}(x^{\prime}_{1},\ldots,x^{\prime}_{n},\zeta)$
-4:  $((c^{\prime}_{1},\dots,c^{\prime}_{k}),cost(S^{\prime}_{\epsilon}))\leftarrow NonPrivateApprox(Y,k)$
-5:  $cost(S_{\epsilon})=\left(\frac{\log(n/\beta)}{0.01}\right)^{p/2}cost(S^{\prime}_{\epsilon})$
-6:  $c\leftarrow\textsc{DimReverse}^{\epsilon/2}((c^{\prime}_{1},\dots,c^{\prime}_{k}),(x^{\prime}_{1},\dots,x^{\prime}_{n}))$
-7:  return $c=\{c_{1},\ldots,c_{k}\},cost(S_{\epsilon})$, $Y$, $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
-Algorithm 4 PrivateExplanations
-Input: $Y,\text{cost}(S_{\epsilon}),V_{s},k,\epsilon,p,n,\beta$
- Output: {$cost(S_{\epsilon})-cost(S^{(i)}_{\epsilon}):i\in V_{s}\}$ DP explanations for each agent (i)
-1:  for $i\in V_{s}$ do
-2:     $cost(S^{(i)}_{\epsilon})\leftarrow\left(\frac{\log(n/\beta)}{0.01}\right)^{\frac{p}{2}}NonPrivateApproxFC(Y,k,x^{\prime}_{i})$
-3:  end for
-4:  return $\{cost(S_{\epsilon})-cost(S^{(i)}_{\epsilon}):i\in V_{s}\}$
-Private contrastive explanation:
-The following summarizes Algorithm 4 - PrivateExplanations:
-Input: Private coreset ($Y$), Private clustering cost ($cost(S_{\epsilon})$), Agents looking for contrastive explanation ($V_{s}$).
-Output: Explanation for each agent in $V_{s}$ ( $\text{cost}(S_{\epsilon})-cost(S^{(i)}_{\epsilon}$)).
-  1. 1.
-    Fixed-centroid clustering on Coreset - NonPrivateApproxFC (specifically for $k$-means and $k$-median in Section 4.2 and Section 4.1 respectively)
-  2. 2.
-    Calculate clustering cost by multiplication - Algorithm 4, line 2.
-  3. 3.
-    Return utility decrease as an explanation - Algorithm 4, line 4.
-After the completion of PrivateClustering, which yields private centroids, specific users within $V_{s}\subset\mathbf{X}$, may request an explanation regarding the absence of a facility at their particular location.
-We want to provide each agent with the overall cost overhead that a facility constraint at his specific location will yield.
-To address these queries, we integrate our peCluster mechanism.
-This algorithm takes the previously computed coreset and employs our NonPrivateApproxFC(tailor-suited non-private algorithm we designed and explained in the following sections) $w^{\prime}$-approximate algorithm to output the cost of $(k,p)$-clustering with a fixed centroid.
-The output of our algorithm provides the difference cost of the clustering, denoted as cost($S^{(i)}_{\epsilon}$) - cost($S_{\epsilon}$).
-This output effectively captures the contrast between the overall clustering cost and the specific clustering cost when an individual agent ($i$) inquires about the absence of a facility at their location, aligning precisely with the key objectives and functions outlined earlier in the abstract.
-The privacy analysis, as demonstrated in Theorem 1, establishes the privacy guarantees of PrivateClustering and peCluster.
-The output $Y$ is $\epsilon/2$-differentially private as an output of $\epsilon/2$-DP algorithm.
-Consequently, both $(c^{\prime}_{1},\ldots,c^{\prime}_{k})$ and $cost(S_{\epsilon})$ maintain $\epsilon/2$-DP status, under the Post-Processing property.
-Applying $\textsc{DimReverse}^{\epsilon/2}$ to find the centers in the original space, $c=\{c_{1},\ldots,c_{k}\}$ is $\epsilon$-DP by Composition theorem.
-For each $i$, cost($S^{(i)}_{\epsilon}$) is produced by Post-Processing of $Y$ with only $x^{\prime}_{i}$, hence $cost(S^{(i)}_{\epsilon}$) satisfies $\epsilon$-$i$-exclusive-DP.
+DimReverse, which includes the FindCenter operation (detailed in the Appendix), has a time complexity of $O(\text{poly}(np))$ and is invoked $k$ times.
+Additionally, we execute one standard $(k,p)$-clustering and $|X|$ instances of $(k,p)$-clustering with a fixed center.
+Together, these steps ensure the algorithm’s overall polynomial runtime.
+All symbols used in this analysis are defined in Table 2 in the Appendix.
 Theorem 2.
-(Full proof in Theorem 5.) The solution $(c_{1},\ldots,c_{k})$ and $cost(S_{\epsilon})$ released by Algorithm 3 are $\epsilon$-DP.
-For all clients $i$, the value $S^{(i)}_{\epsilon}$ released by Algorithm 4 is $\epsilon$-$x_{i}$-exclusive DP.
-A critical component of our algorithm involves a detailed utility analysis. In the following sections, we will present rigorous upper bounds and specific constraints for k-means and k-median.
-Utility Analyses.
-PrivateCoreset is setup with parameter $\zeta$ (other than the privacy budget $\epsilon/2$), which depends on $\alpha$ (Line 1, Algorithm 3).
-It leads to the cost of every clustering on $Y$ (the coreset) is off from its cost on $X^{\prime}$ by a multiplicative factor of $(1+0.1\alpha)$ plus an additive factor of $\overset{\sim}{O}(\text{polylog}(n/\beta)/\epsilon)$.
-Then, by applying the Dimensional Reduction lemma (Appendix section B), which states that the cost of a specific clustering on $X^{\prime}$ ($d^{\prime}$-dimensional space) is under some constant factor of the same clustering on $X$ ($d$-dimensional space), we can bound the $cost({S_{\epsilon}}^{(i)})$ by its optimal clustering $OPT_{i}$.
-Lemma 1.
-(Full proof in Lemma 7) With probability at least $1-\beta$, $S_{\epsilon}$ (clustering cost) is a $(w,t)$-approximation of $OPT$, where [^1] We use the notation $\mathcal{O}_{p,\alpha}$ to explicitly ignore factors of $p,\alpha$:
-$w$ $=w^{\prime\prime}(1+\alpha),$ (1)
-$t$ $=w^{\prime\prime}\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.\text{polylog}(n/\beta)/\epsilon\right)$ (2)
-We note that $\Lambda$ is chosen carefully by composing (1) the factor of $(d/d^{\prime})^{1/2}$, which will scale to $(d/d^{\prime})^{p/2}$ due to .$\|.\|_{p}$ operator and cancel out the scaling factor $(d^{\prime}/d)^{p/2}$ occurs when mapping the cost of clustering from the lower dimensional space to the original space and (2) the factor of $0.01/\log(n/\beta)$ that guarantees each $\|x^{\prime}_{i}\|<1/\Lambda$ with probability at least $1-0.1\beta/n$ (and for all $i$ simultaneously with probability $1-0.1\beta$).
+Assume there exist polynomial-time algorithms for $(k,p)$-clustering and $(k,p)$-clustering with a fixed center.
+The total running time of Algorithm 1 is $O((k/\beta)^{O_{p,\alpha}(1)}\text{poly}(nd))$.
+This computational complexity demonstrates that our algorithm is efficient for large datasets, balancing the additional overhead of fixed-centroid clustering with practical runtimes.
+Theorem 2 follows from the detailed steps, as PrivateCoreset and FindCenter contribute manageable computational overhead.
+Finally, the algorithm integrates a critical utility analysis to ensure robust performance.
+In the following sections, we present rigorous upper bounds and specific constraints for $k$-means and $k$-median, illustrating the practicality and effectiveness of our approach.
+Utility Analysis.
+PrivateCoreset uses parameters $\zeta$ (which is a function of $\alpha$) and privacy budget $\epsilon/2$, derived from Ghazi et al.
+(2020) and detailed in our Appendix.
+This algorithm produces a coreset $Y$ that ensures the clustering cost on $Y$ closely approximates the cost on the projected dataset $X^{\prime}$.
+Specifically, the approximation is within a $(1+0.1\alpha)$ factor, plus an additive $\overset{\sim}{O}(\text{polylog}(n/\beta)/\epsilon)$ term.
+Then, by applying the Dimensional Reduction lemma (in the Appendix), which states that the cost of a specific clustering on $X^{\prime}$ ($d^{\prime}$-dimensional space) is under some constant factor of the same clustering on $X$ ($d$-dimensional space), we can bound the $cost({S_{\epsilon}}^{(i)})$ by its optimal clustering $OPT_{i}$.
+We first state the approximation factor derived using Ghazi et al.
+(2020), since this is used in our analysis.
 Theorem 3.
-(Full proof in Theorem 6) Fix an $i$.
-With probability at least $1-\beta$, $S^{(i)}_{\epsilon}$ released by Algorithm 4 is a $(w,t)$-approximation of $OPT_{i}$-the optimal $(k,p)$-clustering cost with a center fixed at position $x_{i}$, in which:
-$w$ $=w^{\prime}(1+\alpha)$ (3)
-$t$ $=w^{\prime}\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.\text{polylog}(n/\beta)/\epsilon\right)$ (4)
-Lemma 2.
-(Full proof in Lemma 8) Fix an $i$.
-If $OPT_{i}\geq w^{\prime\prime}(1+\alpha)OPT+t^{(i)}$, then with probability at least $1-2\beta$, ${S_{\epsilon}}$ and ${S_{\epsilon}}^{(i)}$ released by Algorithm 3 and 4 satisfies that ${S_{\epsilon}}^{(i)}>{S_{\epsilon}}$.
-Running Time Analysis.
-Assume that polynomial-time algorithms exist for $(k,p)$-clustering and for $(k,p)$-clustering with a fixed center.
-Setting $d^{\prime}=O(p^{4}\log(k/\beta))$ (that satisfies Dimension-Reduction-Lemma(Appendix section B)), algorithm PrivateCoreset takes $O(2^{O(d^{\prime})}\text{poly}(n))$ time (Lemma 42 of Ghazi et al.
-(2020)), which is equivalent to $O((k/\beta)^{O_{p,\alpha}(1)}\text{poly}(n))$.
-FindCenter takes $O(\text{poly}(np))$ time (Corollary 54 of Ghazi et al.
-(2020)) and is being called $k$ times in total.
-Finally, we run $1$ instance of $(k,p)$-clustering and $|V_{s}|$ instances of $(k,p)$-clustering with a fixed center.
-It follows that the total running time of our algorithm is $O((k/\beta)^{O_{p,\alpha}(1)}\text{poly}(nd))$.
-This means the computational complexity of our algorithm is polynomial in relation to the size of the input data.
+Cost of explanations due to privacy. Fix an agent $i$. With probability at least $1-\beta$, cost($S^{(i)}_{\epsilon}$) computed by Algorithm peCluster is a $(w,t)$-approximation of $OPT_{i}$, with
+$w$ $=w^{\prime\prime}(1+\alpha)$
+$t$ $=w^{\prime\prime}\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.\text{polylog}(n/\beta)/\epsilon\right)$
+As $S^{(i)}_{\epsilon}$ results from a randomized mechanism, its cost is higher than $S_{\epsilon}$’s most of the time with high probability, ensuring a positive private explanation.
 Tight Approximation Ratios.
-Corollary 1 and 2 will wrap up this section by presenting a specific, tight approximation ratio($w^{\prime}$) achieved after applying NonPrivateApproxFC.
-These corollaries will detail and confirm the algorithm’s effectiveness in achieving these approximation ratios.
+The most challenging aspect of our analysis is determining the precise approximation factor $w^{\prime\prime}$ for k-means and k-median in the context of fixed-centroid clustering.
+In the following sections, we will present modifications to well-known k-median and k-means algorithms, adapting them for fixed-centroid clustering scenarios.
+We will then demonstrate that these modified algorithms achieve the same tight approximation factors.
+Formally, we show how the well-known utility bounds of k-means and k-median can be preserved while fixing one centroid to a requested location, ensuring the robustness of these algorithms under such constraints.
+Corollary 1 and Corollary 2 will conclude this section by presenting the specific, tight approximation ratios ($w^{\prime\prime}$) achieved after applying our NonPrivateApproxFC algorithm.
+These corollaries will provide detailed confirmation of our algorithm’s effectiveness in achieving these optimal approximation ratios within the constraints of differential privacy.
 4.1 NonPrivateApproxFC for $k$-median
-As previously mentioned, we have specifically developed a non-private fixed centroid clustering algorithm, referred to as NonPrivateApproxFC.
-Our algorithm is proved to be an 8-approximation algorithm.
-We obtain our results by adapting Charikar et al.
-(1999) to work with a fixed centroid (Moving forward in our discussion, we will refer to the fixed centroid as $C$).
-To grasp how we adapted the algorithm to suit our needs, it’s essential to understand the definitions used in  Charikar et al.
-(1999) as they introduce demand $d_{j}$ for each location $j\in N$ as a weight reflecting the location’s importance.
-$N$ represents list of agents ${1\ldots n}$.
+We have developed a non-private fixed centroid clustering algorithm, which we call NonPrivateApproxFC.
+This algorithm is an adaptation of Charikar et al.
+(1999).
+In the following section, we will prove that our modified algorithm, which works with a fixed centroid (referred to as $z$), achieves an 8-approximation factor.
+To grasp how we adapted the algorithm to suit our needs, it’s essential to understand the symbols used in  Charikar et al.
+(1999).
+In this section, we adopt the notation from Charikar et al.
+(1999) to avoid confusion with the symbols used in this paper, where $d$ and $d^{\prime}$ denote the original and reduced dimensions, respectively.
+In their work, $d_{j}$ represents the demand at each location $j\in N$, serving as a weight that reflects the importance of the location.
+$N$ refers to the set of agents ${1,\ldots,n}$.
 For the conventional k-median problem, each $d_{j}$ is initially set to 1 for all $j\in N$.
-The term $c_{ij}$ represents the cost of assigning any $i$ to $j$ and $x_{ij}$ represents if location $j$ is assigned to center $i$.
-We first assume that the fixed center is one of the input data points $N$ to formulate the $k$-median problem with a fixed center as the solution of the integer program (in the Appendix).
-We can then relax the IP condition to get the following LP:
+The term $c_{ij}$ represents the cost of assigning any $i$ to $j$, $x_{ij}$ represents if location $j$ is assigned to center $i$ and $y_{i}$ indicates if the location i is selected as a center.
+We assume the fixed center is one of the input data points $N$.
+Charikar et al.
+(1999) demonstrates that the $k$-median problem can be formulated as an integer programming problem, and in order to adapt the algorithm we add a constraint in line 9 to treat $z$ as a fixed centroid.
+This modification allows the algorithm to account for the fixed centroid requirement.
+We then relax the integer program (IP) into a linear program (LP) and show that it preserves the same utility bound as the original algorithm.
+By specifying that $y_{z}\geq 1$, we ensure that $y_{z}$ is designated as a centroid in our linear programming formulation.
+Throughout the solution process, $y_{z}$ remains fixed as a centroid.
 minimize $\sum_{i,j\in N}d_{j}c_{ij}x_{ij}$ (1)
@@
 $x_{ij},y_{i}$ $\geq 0\text{ for each $i,j\in N$}$ (4)
-$y_{C},x_{CC}$ $\geq 1\text{ for a fixed $C\in N$}$ (9)
+$y_{z},x_{zz}$ $\geq 1\text{ for a fixed $z\in N$}$ (5)
 Let $(\bar{x},\bar{y})$ be a feasible solution of the LP relaxation and let $\bar{C}_{j}=\sum_{i\in N}c_{ij}\bar{x}_{ij}$ for each $j\in N$ as the total (fractional) cost of client $j$.
+Throughout the three steps, we demonstrate that solving this linear program with the added constraint does not introduce any additional approximation factor.
+The program is solved with the same efficiency and accuracy as it would be without the fixed centroid constraint.
 The first step.
 We group nearby locations by their demands without increasing the cost of a feasible solution $(\bar{x},\bar{y})$, such that locations with positive demands are relatively far from each other.
-By re-indexing, we get $\bar{C}_{C}\leq\bar{C}_{1}\leq\bar{C}_{2}\leq\ldots\bar{C}_{n}$.
-We will show that it’s always possible to position $\bar{C}_{C}$ as the first element of the list, i.e., $\bar{C}_{C}$ is equal to the minimum value of all $\bar{C}_{j}$.
-Recall that: $\bar{C}_{C}=\sum_{i\in N}c_{iC}\bar{x}_{iC}=\sum_{i\in N,i\neq C}c_{iC}\bar{x}_{iC}+c_{CC}\bar{x}_{CC}=0,$ since we know that $\sum_{i\in N}x_{iC}=1,x_{CC}\geq 1$ and $c_{CC}=0$.
+By re-indexing, we get $\bar{C}_{z}\leq\bar{C}_{1}\leq\bar{C}_{2}\leq\ldots\bar{C}_{n}$.
+We will show that it’s always possible to position $\bar{C}_{z}$ as the first element of the list, i.e., $\bar{C}_{z}$ is equal to the minimum value of all $\bar{C}_{j}$.
+Recall that: $\bar{C}_{z}=\sum_{i\in N}c_{i{z}}\bar{x}_{i{z}}=\sum_{i\in N,i\neq z}c_{i{z}}\bar{x}_{i{z}}+c_{zz}\bar{x}_{zz}=0,$ since we know that $\sum_{i\in N}x_{i{z}}=1,x_{zz}\geq 1$ and $c_{zz}=0$.
 The remaining work of the first step follows Charikar et al.
@@
 Demand shift occurs as follows: $d^{\prime}_{i}\leftarrow d^{\prime}_{i}+d^{\prime}_{j},d^{\prime}_{j}\leftarrow 0$.
-Since we initialize $d^{\prime}_{C}=d_{C}=1$, and we never move its demands away, it follows that $d^{\prime}_{C}>0$.
+Since we initialize ${d^{\prime}_{z}}=d_{z}=1$, and we never move its demands away, it follows that $d^{\prime}_{z}>0$.
 Let $N^{\prime}$ be the set of locations with positive demands $N^{\prime}=\{j\in N,d^{\prime}_{j}>0\}$.
 A feasible solution to the original demands is also a feasible solution to the modified demands.
-By Lemma 11, for any pair $i,j\in N^{\prime}$: $c_{ij}>4\max(\bar{C}_{i},\bar{C}_{j})$.
+Lemma 3.
+Locations $i,j\in N^{\prime}$ satisfy: $c_{ij}>4\max(\bar{C}_{i},\bar{C}_{j})$.
+Proof.
+The lemma follows the demands moving step (in the first step of the algorithm): for every $j$ to the right of $i$ (which means $\bar{C}_{j}\geq\bar{C}_{i}$) and within the distance of $\bar{C}_{j}$ (that also covers all points within distance $\bar{C}_{i}$), we move all demands of $j$ to $i$, hence $j$ will not appear in $N^{\prime}$.
+∎
 Lemma 4.
@@
 We will group fractional centers from the solution $(\bar{x},\bar{y})$ to create a new solution $(x^{\prime},y^{\prime})$ with cost at most $2\bar{C}_{LP}$ such that $y^{\prime}_{i}=0$ for each $i\notin N^{\prime}$ and $y^{\prime}_{i}\geq 1/2$ for each $i\in N^{\prime}$.
-We also ensure that $y^{\prime}_{C}\geq 1/2$ in this step, i.e., $C$ will be a fractional center after this.
-The next lemma states that for each $j\in N$, there are at least $1/2$ of the total fractional centers within the distance of $2\bar{C}_{j}$.
+We also ensure that $y^{\prime}_{z}\geq 1/2$ in this step, i.e., ${z}$ will be a fractional center after this.
+A solution is called $1/2$-restricted if $y_{j}\geq 1/2$ for any point $j\in N$ and $y_{j}=0$ otherwise.
+This restriction balances the assignment of demand, ensuring that no single center dominates excessively.
+The concept of $1/2$-restricted solutions is used to create more equitable distributions of demand and is key to transitioning to a $\{1/2,1\}$-integral solution.
+The next lemma leverages this property:
 Lemma 5.
@@
 We now construct a $\{1/2,1\}$-integral solution $(\hat{x},\hat{y})$ with no greater cost.
-Sort the location $j\in N^{\prime},j\neq C$ in the decreasing order of the weight $d^{\prime}_{j}c_{c_{s}(j)j}$ and put $C$ to the first of the sequence, set $\hat{y}_{j}=1$ for the first $2k-n^{\prime}$ locations and $\hat{y}_{j}=1/2$ for the rest.
-By doing that, we minimize the cost by assigning heaviest weights $d^{\prime}_{j}c_{c_{s}(j)j}$ to the maximum multiplier (i.e., $1$) while assigning lightest weights $d^{\prime}_{j}c_{c_{s}(j)j}$ to the minimum multiplier (i.e., $1/2$) for each $j\in N^{\prime},j\neq C$.
-Any feasible $1/2$-restricted solution must have $y^{\prime}_{C}=1$ to satisfy the constraint of $C$ so that the contribution of $\hat{y}_{C}$ is the same as its of $y^{\prime}_{C}$.
+Sort the location $j\in N^{\prime},j\neq{z}$ in the decreasing order of the weight $d^{\prime}_{j}c_{s(j)j}$ and put ${z}$ to the first of the sequence, set $\hat{y}_{j}=1$ for the first $2k-n^{\prime}$ locations and $\hat{y}_{j}=1/2$ for the rest.
+By doing that, we minimize the cost by assigning heaviest weights $d^{\prime}_{j}c_{s(j)j}$ to the maximum multiplier (i.e., $1$) while assigning lightest weights $d^{\prime}_{j}c_{s(j)j}$ to the minimum multiplier (i.e., $1/2$) for each $j\in N^{\prime},j\neq z$.
+Any feasible $1/2$-restricted solution must have $y^{\prime}_{z}=1$ to satisfy the constraint of ${z}$ so that the contribution of $\hat{y}_{z}$ is the same as its of $y^{\prime}_{z}$.
 It follows that the cost of $(\hat{x},\hat{y})$ is no more than the cost of $(x^{\prime},y^{\prime})$.
@@
 (1999) that converts a $\{1/2,1\}$-integral solution to an integral solution with the cost increases at most by $2$.
-We note that there are two types of center $\hat{y}_{j}=1/2$ and $\hat{y}=1$, hence there are two different processes.
+We note that there are two types of center $\hat{y}_{j}=1/2$ and $\hat{y}_{j}=1$, hence there are two different processes.
 All centers $j$ with $\hat{y}_{j}=1$ are kept while more than half of centers $j$ with $\hat{y}_{j}=1/2$ are removed.
-Since we show that $\hat{y}_{C}=1$ in the previous step, $C$ is always chosen by this step and hence guarantees the constraint of $C$.
+Since we show that $\hat{y}_{z}=1$ in the previous step, $z$ is always chosen by this step and hence guarantees the constraint of $z$.
 Theorem 4.
-For the metric k-median problem, the algorithm above outputs an $8$-approximation solution.
+Approximation factor of fixed centroid k-median. For the metric k-median problem, the algorithm above outputs an $8$-approximation solution.
 Proof.
 It is obvious that the optimal of the LP relaxation is the lower bound of the optimal of the integer program.
-While constructing an integer solution for the LP relaxation with the modified demands, Lemma 13 incurs a cost of $2\bar{C}_{LP}$ and the third step multiplies the cost by a factor of $2$, making the cost of the solution (to the LP) at most $4\bar{C}_{LP}$.
-Transforming the integer solution of the modidied demands to a solution of the original input adds an additive cost of $4\bar{C}_{LP}$by Lemma 12 and the Theorem follows.
+While constructing an integer solution for the LP relaxation with the modified demands,  Charikar et al.
+(1999) states that there is a 1/2-restricted solution ($x^{\prime}$, $y^{\prime}$) which costs at most $2\bar{C}_{LP}$.
+And now the third step multiplies this cost by a factor of $2$, making the cost of the solution (to the LP) at most $4\bar{C}_{LP}$.
+Transforming the integer solution of the modified demands to a solution of the original input adds an additive cost of $4\bar{C}_{LP}$ by Lemma 4 of Charikar et al.
+(1999) and the Theorem follows.
 ∎
+Having demonstrated that our modification of  Charikar et al.
+(1999) to execute fixed-centroid k-median instead of standard k-median yields an 8-factor approximation of the optimal solution, we can now proceed to prove that our private explanation closely approximates the optimal solution for the fixed-centroid scenario.
 Corollary 1.
-Running peCluster with NonPrivateApproxFC be the above K-median algorithm, with probability at least $1-\beta$, $S^{(i)}_{\epsilon}$ is a $(w,t)$-approximation of $OPT_{i}$–the optimal K-median with a center fixed at position $x_{i}$, in which:
+Running peCluster with NonPrivateApproxFC be the above K-median algorithm, with probability at least $1-\beta$, $S^{(i)}_{\epsilon}$ is a $(w,t)$-approximation of $OPT_{i}$–the optimal K-median with a center fixed at position $z_{i}$, in which:
 $w$ $=8(1+\alpha)$
@@
 4.2 NonPrivateApproxFC for $k$-means
-In this section, we describe an algorithm that is a $25$-approximation of the k-means problem with a fixed center.
+In this section, we present our NonPrivateApproxFC algorithm for k-means with a fixed center.
+Based on  Kanungo et al.
+(2002), we achieve a 25-approximation.
+We will analyze this approximation factor in detail below.
 We adapt the work by Kanungo et al.
 (2002) by adding a fixed center constraint to the single-swap heuristic algorithm.
-Similar to Kanungo et al.
-(2002), we need to assume that we are given a discrete set of candidate centers $C$ from which we choose $k$ centers.
+As in their result, we need to assume that we are given a discrete set of candidate centers $C$ from which we choose $k$ centers.
 The optimality is defined in the space of all feasible solutions in $C$, i.e., over all subsets of size $k$ of $C$.
-We then present how to remove this assumption, with the cost of a small constant additive factor, in the Appendix.
+We then present how to remove this assumption, with the cost of a small constant additive factor.
 Definition 7.
-Let $O=(O_{1},O_{2},\ldots,O_{k})$ be the optimal clustering with $O_{1}$ be the cluster with the fixed center $\sigma$.
-A set $C\subset\mathbb{R}^{d}$ is a $\gamma$-approximate candidate center set if there exists $\sigma\in\{c_{1},c_{2},\ldots,c_{k}\}\subseteq C$, such that:
-$cost(c_{1},c_{2},\ldots,c_{k})$ $\leq(1+\gamma)cost(O),$ (11)
-Given $u,v\in\mathbb{R}^{d}$, let $\Delta(u,v)$ denote the squared Euclidean distance between $u,v:\Delta(u,v)=dist^{2}(u,v)$.
-Let $S\subset\mathbb{R}^{d}$, $\Delta(S,v)=\sum_{u\in S}\Delta(u,v)$.
-Let $P\subset\mathbb{R}^{d}$, $\Delta_{P}(S)=\sum_{q\in P}\Delta(q,s_{q})$, where $s_{q}$ is the closest point of $S$ to $q$.
-We drop $P$ when the context is clear.
-Let $\sigma$ be the fixed center that must be in the output.
-Let $C$ be the set of candidate centers, that $\sigma\in C$.
-We define stability in the context of $k$-means with a fixed center $\sigma$ as follows.
+Let $O=(O_{1},O_{2},\ldots,O_{k})$ be the optimal clustering with $O_{1}$ be the cluster with the fixed center $z$.
+A set $C\subset\mathbb{R}^{d}$ is a $\gamma$-approximate candidate center set if there exists $z\in\{c_{1},c_{2},\ldots,c_{k}\}\subseteq C$, such that: $cost(c_{1},c_{2},\ldots,c_{k})\leq(1+\gamma)cost(O).$
+Given $u,v\in\mathbb{R}^{d}$, let $\Delta(u,v)$ denote the squared Euclidean distance between $u$ and $v$: $\Delta(u,v)=\text{dist}^{2}(u,v)$.
+For a set $S\subset\mathbb{R}^{d}$, the total squared distance between all points in $S$ and a point $v$ is given by $\Delta(S,v)=\sum_{u\in S}\Delta(u,v)$.
+Similarly, for a set $P\subset\mathbb{R}^{d}$, $\Delta_{P}(S)$ represents the total squared distance between each point $q\in P$ and its closest point $s_{q}\in S$.
+Here, $q$ refers to an individual data point in set $P$, and $s_{q}$ is its nearest neighbor in $S$.
+When the context is clear, we drop $P$ for simplicity.
+This notation captures the essential relationships between points and their nearest centroids.
+Let $z$ be the fixed center that must be in the output.
+Let $C$ be the set of candidate centers, that $z\in C$.
+We define stability in the context of $k$-means with a fixed center $z$ as follows.
 We note that it differs from the definition of Kanungo et al.
-(2002) such that we never swap out the fixed center $\sigma$:
+(2002) such that we never swap out the fixed center $z$:
 Definition 8.
-A set $S$ of k centers that contains the fixed center $\sigma$ is called $1$-stable if: $\Delta\big{(}S\setminus\{s\}\cup\{o\}\big{)}\geq\Delta(S),$ for all $s\in S\setminus\{\sigma\}$, $o\in O\setminus\{\sigma\}$.
-Algorithm. We initialize $S^{(0)}$ as a set of $k$ centers form $C$ that $\sigma\in S^{(0)}$. For each set $S^{(i)}$, we perform the swapping iteration:
-  • Select one center $s\in S^{(i)}\setminus\sigma$
+A set $S$ of k centers that contains the fixed center $z$ is called $1$-stable if: $\Delta\big{(}S\setminus\{s\}\cup\{o\}\big{)}\geq\Delta(S),$ for all $s\in S\setminus\{z\}$, $o\in O\setminus\{z\}$.
+Algorithm. We initialize $S^{(0)}$ as a set of $k$ centers form $C$ that $z\in S^{(0)}$. For each set $S^{(i)}$, we perform the swapping iteration:
+  • Select one center $s\in S^{(i)}\setminus z$
   • Select one replaced center $s^{\prime}\in C\setminus S^{(i)}$
@@
 Theorem 5.
-If $S$ is an $1$-stable $k$-element set of centers, then $\Delta(S)\leq 25\Delta(O)$.
-Furthermore, if $C$ is a $\frac{\gamma}{25}$-approximate candidate center set, then $S$ is a $(25+\gamma)$-approximate of the actual optimal centroids in the Euclidean space.
-Kanungo et al.
-(2002)’s analysis depends on the fact that all centers in the optimal solution are centroidal, hence they can apply centroidal-Lemma in the Appendix.
-Our analysis must take into account that $\sigma\in O$ is a fixed center, which is not necessarily centroidal.
-Therefore, we need to adapt and give special treatment for $\sigma$ as an element of the optimal solution.
-Also, in the swapping iteration above, we never swap out $\sigma$, since the desired output must always contain $\sigma$, which also alters the swap pairs mapping as in Kanungo et al.
-(2002).
-We adapt the mapping scheme of Kanungo et al.
-(2002), described in the Appendix with one modification: we set up a fixed capturing pair between $\sigma\in S$ and $\sigma\in O$.
-It makes sense since $\sigma=o_{\sigma}=s_{\sigma}$, as the closest center in $S$ to $\sigma$ is itself.
-With it, we always generate a swap partition of $S_{j}$ and $O_{j}$ that both contain $\sigma$.
-The reason we must have this swap partition (even though we never actually perform this swap on $\sigma$) is that we will sum up the cost occurred by every center in $S$ and $O$ by pair-by-pair from which the total costs $\Delta(S)$ and $\Delta(O)$ are formed (in Lemma 5).
-For this partition, there are two cases:
-  • $|S_{j}|=|O_{j}|=1$, which means the partitions contain only $\sigma$, and hence when they are swapped, nothing changes from $S$ to $S^{\prime}$ and hence do not violate the constraint we set above about not swapping out $\sigma$.
-  • $|S_{j}|=|O_{j}|>1$, which means the partitions contain some $o\in O$ that $s_{o}=\sigma$.
-By the design of the swap pairs mapping, all $s\in S_{j}\setminus\sigma$ will be chosen to be swapped out instead, and $\sigma\in S_{j}$ will never be swapped out.
-For each swap pair of $s$ and $o$, we analyze the change in the distortion.
-For the insertion of $o$, all points in $N_{O}(o)$ are assigned to $o$, leads to the change of distortion to be: $\sum_{q\in N_{O}(o)}\big{(}\Delta(q,o)-\Delta(q,s_{q})\big{)}$.
-For all points in $N_{S}(s)$ that lost $s$, they are assigned to new centers, which changes the distortion to be:
-$\sum_{q\in N_{S}(s)\setminus N_{O}(o)}$ $\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)}\leq\sum_{q\in N_{S}(s)}\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)},$
-and by the property $3$ of swap pairs, $s_{o_{q}}\neq s$, since $o_{q}\neq o$; the inequality is because for all $q$, $\Delta(q,s_{o_{q}})-\Delta(q,s)\geq 0$.
-Lemma 5.
-(Full proof in Lemma 15) Let $S$ be $1$-stable set and $O$ be the optimal set of $k$ centers, we have $\Delta(O)-3\Delta(S)+2R\geq 0$, where $R=\sum_{q\in P}\Delta(q,s_{o_{q}})$.
-Lemma 6.
-(Full proof in Lemma 17) With $R$ and $\alpha$ defined as above: $R\leq 2\Delta(O)+(1+2/\alpha)\Delta(S)$.
-Proof of Theorem 4. By the results of Lemma 5 and 6:
-00 0 $\leq\Delta(O)-3\Delta(S)+2R$ (12)
-$\leq\Delta(O)-3\Delta(S)+4\Delta(O)+2\Delta(S)+\frac{4}{\alpha}\Delta(S)$ (13)
-$=5\Delta(O)-(1-4/\alpha)\Delta(S).$ (14)
-Using the same argument with Kanungo et al.
-(2002), we have $5/(1-4/\alpha)\geq\alpha^{2}$ or $(\alpha-5)(\alpha+1)\geq 0$ or $\alpha\leq 5$ which means $\frac{\Delta(S)}{\Delta(O)}=\alpha^{2}\leq 25$, and the Theorem follows.
-In Section D.2, we describe how to create a $\gamma$-approximation candidate center set for any constant $\gamma$.
-Using it to create $C$ as a $\frac{25}{\gamma}$-approximation candidate center set and perform the above algorithm on $C$, the final output $S$ will be a $(25+\gamma)$-approximation.
+Approximation factor of fixed centroid k-mean.
+If $S$ is an $1$-stable $k$-element set of centers, $\Delta(S)\leq 25\Delta(O)$.
+Furthermore, if $C$ is a $\frac{\gamma}{25}$-approximate candidate center set, $S$ is a $(25+\gamma)$-approximate of the actual optimal centroids in the Euclidean space.
+Having demonstrated that our modification of Kanungo et al.
+(2002) to execute fixed-centroid k-means instead of standard k-means yields a 25-factor approximation of the optimal solution, we can now proceed to prove that our private explanation closely approximates the optimal solution for the fixed-centroid scenario.
 Corollary 2.
@@
 $t$ $=(25+\gamma)\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.polylog(n/\beta)/\epsilon\right).$
-Figure 1: A detailed visualization of our dataset (Albemarle County, Virginia) and analysis includes (a) a Scatter plot of the full dataset with sampled points for contrastive analysis.
+With the utility bounds for k-means and k-median under the fixed-centroid constraint proven, it is clear that altering the original algorithms preserves the same utility bounds as their non-fixed counterparts.
+This ensures that accommodating fixed centroids does not compromise clustering quality.
+Notably, these bounds refer to clustering utility, not explanation bounds.
+By showing that fixed-k-means and fixed-k-median perform as effectively as standard versions, users can trust the quality of the explanations.
+Without these guarantees, users might question the validity of centroid placements.
+Our results ensure the explanations are based on clustering solutions that are as robust and reliable as the original algorithms.
+Figure 1: A visualization of our dataset (Heart Disease dataset from the UCI ML Repository), projected into an 8-dimensional space.
+(a) t-SNE of our data.
 (b) Comparison of k-median clustering with fixed and non-fixed centroids, both private and non-private.
 (c) Bar graph showing contrastive explanation differences for differential private and non-private k-median with a fixed centroid.
-(d) Heatmap indicating k-median error costs based on fixed centroid choices.
+(d) We fix the privacy budget of 0.6 while demonstrating the contrastive explanation across various dimensions.
 5 Experiments
-In this part of our study, we focus on understanding the impact of the privacy budget $\epsilon$ on the trade-off between privacy and accuracy.
-Additionally, we explore how $\epsilon$ influences the quality of differentially private explanations.
-To assess the impact of $\epsilon$ on privacy and accuracy, we will employ a set of metrics: 1) PO (Private Optimal) to evaluate the clustering cost using the differentially private algorithm, 2) PC (Private Contrastive) to determine the cost in fixed centroid differentially private clustering, 3) RO (Regular Optimal) to measure clustering cost in regular, non-private data, and 4) RC (Regular Contrastive) to analyze the cost of fixed centroid clustering in non-private data.
-Measurements To measure the quality of explanations, we define two key metrics: 1) APE (Average Private Explanation), which calculates the difference between PC and PO to represent the decrease in utility for clustering as an explanatory output for agents, and 2) AE (Average Explanation), which is the difference between RO and RC, serving as a baseline to assess the extent to which differential privacy impacts the explanations.
-These metrics will enable us to effectively quantify the explanatory power of our approach.
-We study the trade-offs between privacy and utility for the above metrics. The $\epsilon$ value balances privacy and utility: lower $\epsilon$ values prioritize privacy but may compromise utility.
-Datasets Our study focuses on activity-based population datasets from Charlottesville City and Albemarle County, Virginia, previously used for mobile vaccine clinic deployment Mehrab et al.
+Our study examines how the privacy budget $\epsilon$ affects the trade-off between privacy and accuracy, focusing on the quality of differentially private explanations.
+We use four key metrics: Private Optimal (PO, $S_{\epsilon}$), Private Contrastive (PC, $S^{(i)}_{\epsilon}$), Regular Optimal (RO, $OPT$), and Regular Contrastive (RC, $OPT_{i}$), to compare clustering costs with and without fixed centroids in both private and non-private algorithms.
+To assess explanation quality, we define two derived metrics: Average Private Explanation (APE, PC - PO, cost($S^{(i)}_{\epsilon}$) - cost($S_{\epsilon}$)) and Average Explanation (AE, RC - RO, cost($S^{(i)}$) - cost($S$)).
+APE measures utility loss in private clustering as an explanatory output, while AE provides a non-private baseline.
+These metrics help us evaluate the explanatory power of our approach.
+By analyzing these metrics across different $\epsilon$ values, we explore the balance between privacy and utility, highlighting the trade-offs in our differentially private clustering and explanation framework.
+Datasets Our research utilizes a diverse set of datasets to demonstrate the versatility and effectiveness of our approach, as summarized in Table 1.
+We employed the Heart Disease dataset featuring 13 dimensions, and the Breast Cancer dataset with 30 features, including both numeric and categorical fields.
+Both datasets were taken from the UCI Machine Learning Repository Those higher-dimensional datasets were crucial in validating our theoretical framework.
+Additionally, we used two-dimensional activity-based population datasets from Charlottesville City and Albemarle County, Virginia, previously employed in mobile vaccine clinic deployment studies Mehrab et al.
 (2022).
-The Charlottesville dataset includes 33,000 individuals, and Albemarle County has about 74,000.
-Due to our algorithm’s intensive computational demands, we sampled 100 points for contrastive explanations.
-Computational intensity of our algorithm stems primarily from the k-median algorithm we employ, which formulates the problem using linear programming.
-When the input size is large, processing times for this specific k-median algorithm can increase significantly.
-However, the implementation we use is one of the more straightforward approaches to executing k-median with a fixed centroid, and faster algorithms could potentially be applied to alleviate these concerns.
-For this study, we have opted not to focus on optimizing the computation time but rather to sample the data points to expedite our results.
-Additionally, we chose to sample 100 points because this number strikes a balance between computational efficiency and the clarity of the results for the reader.
-Importantly, the number of sampled points does not substantially impact the outcome of our analysis, allowing us to maintain the integrity of our findings while ensuring readability and accessibility for our audience.
-Furthermore, our method is data agnostic and does not rely on the s-sparse feature for sample points.
-It is sufficiently adaptable to handle sample points from any distribution.
-This flexibility ensures that our approach is robust and versatile, capable of effectively managing diverse datasets regardless of their sparsity or the specific characteristics of their distribution.
-Dimensiality It is important to note that our experiments were conducted on a 2D dataset.
-This choice was driven by the fact that the initial step of our method involves reducing dimensionality, as outlined in DimReduction (Algorithm 1).
-This process entails scaling down and normalizing the data to produce a lower-dimensional dataset.
-By implementing this step at the outset, we effectively avoid the challenges associated with managing high-dimensional datasets, thus simplifying the data preparation phase.
-Experimental results.
-In this section, we focus solely on the Albemarle County dataset, with additional analyses and datasets presented in the Appendix.
-Figure 1(a) displays the dataset, highlighting the original data points with blue dots and the sampled points used for contrastive explanations with orange ’X’ markers.
-In Figure 1(b), we observe significant trends in the k-median clustering results for both private and non-private settings.
-Both the Private Optimal (PO) and Private Contrastive (PC) exhibit a downward trend in the private setting, indicating that as the $\epsilon$ budget increases, the error cost decreases, which underscores the trade-off between privacy and accuracy.
-The trends for non-private settings are depicted as straight lines, reflecting their independence from the privacy budget and are included for comparative purposes.
-Our analysis shows a consistent gap between the Private Optimal (PO) and Private Contrastive (PC) across different levels of $\epsilon$, marking a stable contrastive error that does not fluctuate with changes in the privacy budget.
-This stability is pivotal as it supports our goal of providing dependable explanations to agents, regardless of the $\epsilon$ value, ensuring that our approach adeptly balances privacy considerations with the demand for meaningful explainability.
-The distinction between PO and PC arises because PO is derived from implementing the optimal k-median/k-means algorithm, while PC results from enforcing a centroid’s placement at a specific agent’s location, inherently producing a less optimal outcome due to this constraint.
-Initially, agents are presented with the optimal outcome (PO).
-When an agent inquires why a centroid is not positioned at their specific location, we offer a contrastive explanation (PC), which is naturally less optimal than the original result as it evaluates the utility change from hypothetically placing a centroid at that specific location.
-In this scenario, a higher score signifies lower utility, emphasizing the balance between providing personalized explanations and adhering to privacy constraints.
-Figure 1(c) displays the average private explanation (APE) and average explanation (AE) with variance across different explanations.
-The average values of APE across multiple runs are relatively stable, irrespective of the privacy budget, reinforcing the robustness of our methodology against the constraints imposed by differential privacy.
-Lastly, the heatmap in Figure 1(d) provides insights into how the relative spatial positioning affects the contrastive explanation.
-It vividly shows that the greater the distance a fixed centroid is from the distribution center, the larger the contrastive error, further illustrating the intricate dynamics of our clustering approach.
-6 Discussion
+To complement these real-world datasets, we also generated a synthetic two-dimensional dataset.
+By testing our method on both high-dimensional and two-dimensional data, as well as on real and synthetic datasets, we showcase its robustness across different data complexities and origins.
+Dataset Dim Size Source
+Heart Disease 13 303 UCI MLR
+Breast Cancer Wisconsin 31 569 UCI MLR
+Charlottesville 2 33K Mehrab et al. (2022)
+Albemarle 2 74K Mehrab et al. (2022)
+Synthetic dataset 2 1k Generated
+Table 1: Datasets used in our research
+Figure 2: A visualization of another dataset (Breast Cancer dataset (30 features) from the UCI ML Repository), projected into a 16-dimensional space.
+(a) t-SNE of our data.
+(b) Comparison of k-means clustering with fixed and non-fixed centroids, both private and non-private.
+(c) Bar graph showing contrastive explanation differences for differential private and non-private k-means with a fixed centroid.
+Data Preprocessing: We normalized all datasets to fit within a unit ball to ensure consistency with prior work and standardize our analysis framework.
+While this normalization alters the absolute scale, it preserves the relative relationships between data points, which is crucial for clustering.
+The entire preprocessed dataset was used for analysis, as there is no ground truth labeling for a traditional train-test split in this unsupervised task.
+Dimensionality Our study explored both 2D and higher-dimensional datasets.
+A crucial aspect of our methodology, DimReduction, employs Principal Component Analysis (PCA) for initial dimensionality reduction.
+This process normalizes the data and creates lower-dimensional representations.
+We performed extensive experiments, reducing high-dimensional datasets to various lower dimensions, including 2D, with additional low-dimension experiments.
+Remarkably, our results remained consistent across different reduced dimensionalities.
+Even when reducing data from 13 dimensions to 2, we observed similar trends and results as with other dimensional reductions, despite significant information loss.
+This consistency underscores our method’s robustness across varying dimensions.
+By addressing high-dimensional data challenges through PCA reduction, we ensure our technique’s applicability and efficiency across diverse dataset complexities, maintaining result integrity regardless of original data dimensionality.
+Running Time Analysis: The computational complexity of our algorithm varies by clustering method.
+For k-means, we use the linear-time algorithm from  Kanungo et al.
+(2002), while the k-median approach relies on polynomial-time Linear Programming (LP).
+We have optimized performance with GPU parallelization, reducing execution times from minutes to seconds for both differentially private coresets and clustering tasks.
+Our method is data-agnostic, handling any data distribution efficiently, independent of sparsity.
+For reproducibility, we provide our code, experimental details, and pre-processed datasets in a public repository.
+5.1 Experimental results
+Figure 1 presents four key visualizations of our differentially private clustering and explanation framework across various dimensions.
+The t-SNE plot (leftmost) shows the 2D representation of our dataset, revealing potential clusters and patterns.
+The second plot illustrates K-median cost versus $\epsilon$ for our four metrics (PC, PO, RC, RO), demonstrating the privacy-accuracy trade-off and the consistency of our contrastive explanations.
+In plots (b) and (c) of Figure 1, the x-axis represents the privacy budget $\epsilon$, which we tested over the range $[0,1]$ in intervals of 0.05.
+This granularity allows for a detailed analysis of the privacy-utility trade-off.
+Smaller values of $\epsilon$ enforce stronger privacy guarantees, as reflected in higher clustering costs for PC and PO.
+Conversely, as $\epsilon$ increases, these costs gradually decrease, highlighting the improved utility that comes with relaxed privacy constraints.
+Importantly, the observed trends demonstrate that the framework consistently balances privacy and utility, even at stricter privacy levels.
+Additionally, the stability of the non-private metrics (RC and RO) across the $\epsilon$ range provides a robust baseline for evaluating the performance of our private clustering and explanation methods.
+As expected, both PC and PO costs decrease as $\epsilon$ increases, demonstrating the trade-off between privacy and accuracy.
+The non-private metrics (RC and RO) remain constant across $\epsilon$ values, serving as baselines for comparison.
+Notably, the gap between PC and PO remains relatively consistent, indicating that our contrastive explanations maintain their relative quality at different levels of privacy.
+The third plot illustrates the Explanation Utility for both private (Average Private Explanation) and non-private (Average Non-Private Explanation) scenarios across various $\epsilon$ values.
+This graph quantifies the difference in clustering cost between the optimal solution and the solution with a fixed centroid, representing our contrastive explanations.
+Notably, we observe that the Average Private Explanation remains relatively stable across different $\epsilon$ values.
+This stability is crucial as it indicates that the quality of our contrastive explanations in the private setting is consistent, regardless of the privacy budget.
+The consistent performance across different $\epsilon$ values underscores the robustness of our method, providing reliable explanations even under strict privacy constraints.
+The rightmost plot demonstrates the difference between PC and PO across dimensions for a fixed $\epsilon$, illustrating our method’s scalability with dimensionality.
+Figure 2 follows a similar format but uses another high-dimensional dataset.
+This dataset was reduced from 30 dimensions to 16 to test the robustness of our approach on different datasets and higher dimensions.
+Unlike Figure 1, this figure presents results obtained using the k-means algorithm.
+Furthermore, we extended our experiments to include other reduced dimensions for both k-means and k-median, with the detailed results provided in the Appendix.
+These additional experiments further validate the adaptability and robustness of our framework across different clustering methods and dimensionality settings.
+6 Conclusions
 Our work explores the design of private explanations for clustering, particularly focusing on the k-median and k-means objectives for Euclidean datasets.
-We formalize this as the PrivEC problem, which gives a contrastive explanation to each agent equal to the increase in utility when a cluster centroid is strategically positioned near them.
-The accuracy bounds of our algorithm match private clustering methods, even as they offer explanations for every user, within a defined privacy budget.
-The related work in this domain has seen the development of algorithms for contrastive explanations, but our contribution stands out by integrating differential privacy guarantees.
+We formalize this as the PRIVEC problem, where each agent receives a contrastive explanation corresponding to the loss in utility they experience when a cluster centroid is placed at a strategic position chosen by the agent.
+Our algorithm provides explanations to each user while maintaining the same approximation factor as private clustering, within a predefined privacy budget.
+The related work in this domain has shown the development of algorithms for contrastive explanations, but our contribution stands out by integrating differential privacy guarantees.
 Our experiments demonstrate the resilience of our approach.
-Despite the added layer of providing differential private explanations on top of differential private clustering, the quality of our explanations remains uncompromised.
+Despite the added layer of providing differentially private explanations on top of differentially private clustering, the quality of our explanations remains uncompromised.
+The extended experiments on all our datasets further validate our approach’s efficacy.
+The balance between privacy and utility, the robustness of contrastive explanations, and the negligible impact of $\epsilon$ on explainability were consistent across datasets.
+These findings underscore the potential of our method for diverse real-world applications.
+Our approach is not restricted to k-means and k-median but can be applied to other clustering algorithms as well.
+The methodology leverages fundamental principles common to many clustering techniques, such as centroids and utility functions.
+As long as a clustering algorithm defines centroids and evaluates clustering quality using these metrics, our approach can be adapted to provide privacy-preserving contrastive explanations.
+This adaptability makes it suitable for extending to other paradigms, such as density-based or hierarchical clustering, extending its applicability to various datasets and contexts.
+Acknowledgment This research is partially supported by the Israel Ministry of Innovation, Science & Technology grant 1001818511, NSF grants CCF-1918656, CNS-2317193, IIS-2331315, and CDC MIND U01CK000589.
 Appendix A Related work: additional details
@@
 They showed that their mechanism restricts the information that the adversary can learn about the top important features while maintaining the faithfulness of the explanations.
-The analysis presented in Goethals et al.
+The analysis presented in restatableGoethals et al.
 [2022] considers security concerning contrastive explanations.
@@
 Consequently, all subsequent work, including ours, assumes points are restricted to a unit ball.
+In addition, there has been extensive work on a closely related problem, in the context of private clustering on graphs or networked data, often mentioned as community detection Nguyen and Vullikanti [2024], Chen et al.
+[2023], Ji et al.
+[2019].
+Algorithm 2 DimReduction
+Input: $(x_{1},x_{2},\ldots,x_{n}),d,d^{\prime},\beta$
+Output: $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$ low-dimensional space dataset.
+1:  $\Lambda=\sqrt{\frac{0.01d}{\log(n/\beta)d^{\prime}}}$
+2:  for $i\in\{1,..,n\}$ do
+3:     $\tilde{x}_{i}\leftarrow\Pi_{\mathcal{S}}(x_{i})$
+4:     if $\|\tilde{x}_{i}\|\leq 1/\Lambda$ then
+5:        $x^{\prime}_{i}=\Lambda\tilde{x}_{i}$
+6:     else
+7:        $x^{\prime}_{i}=0$
+8:     end if
+9:  end for
+10:  return $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
+Algorithm 3 DimReverse
+Input: $(c^{\prime}_{1},\ldots,c^{\prime}_{k}),(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
+Output: $(c_{1},\ldots,c_{k})$ Private Centroids in high dimension
+1:  $\mathbf{X}_{1},\ldots,\mathbf{X}_{k}\leftarrow\text{ the partition induced by }$ $(c^{\prime}_{1},\ldots,c^{\prime}_{k})$ on $(x^{\prime}_{1},\ldots,x^{\prime}_{n})$
+2:  for $j\in\{1,\ldots,k\}$ do
+3:     $c_{j}\leftarrow\textsc{FindCenter}^{\epsilon/2}(\mathbf{X}_{j})$
+4:  end for
+5:  return $(c_{1},\ldots,c_{k})$
 Appendix B Additional proofs for peCluster
-Theorem 5.
-(Full version of Theorem 1) The solution $c$ and $\text{cost}{(S_{\epsilon})}$ released by Algorithm 3 are $\epsilon$-DP.
-For all clients $i$, the value $\text{cost}(S_{\epsilon})-\text{cost}{S^{(i)}_{\epsilon}}$ released by Algorithm 4 is $\epsilon$-$x_{i}$-exclusive DP.
+See 1
 Proof.
-It follows that $\text{cost}(S_{\epsilon})$ is the direct results (using no private information) of $Y$, which is $\epsilon/2$-differentially private coreset.
+It follows that $\text{cost}(S_{\epsilon})$ is the direct results of $Y$, which is $\epsilon/2$-differentially private coreset.
 By the post-processing property, $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon$-DP).
-$c$ is the output of DimReverse that is $\epsilon/2$-DP w.r.t.
-the input $(X_{1},X_{2},\ldots,X_{k})$.
-By composition, $c$ is $\epsilon$-DP since the input $(X_{1},X_{2},\ldots,X_{k})$ is calculated partially from $Y$, which is $\epsilon/2$-DP.
-For each explanations $S^{(i)}_{\epsilon}$, let $D,D^{\prime}:D\setminus\{x_{i}\}\sim D^{\prime}\setminus\{x_{i}\}$, let $S^{(i)}_{\epsilon}(Y^{D},x_{i})$ be the value of $S^{(i)}_{\epsilon}$ with input dataset $D$ (and $Y^{D}$ as the private coreset of $D$ respectively) and a fixed center $x_{i}$, let $T=\{Y:S^{(i)}_{\epsilon}(Y,x_{i})\in S\}$, we have $D\sim D^{\prime}$ and therefore:
-$\Pr[S^{(i)}_{\epsilon}(D,x_{i})\in S]$ $=\Pr[Y^{D}\in T]$ (15)
-$\leq e^{\epsilon/2}\Pr[Y^{D^{\prime}}\in T]$ (16)
-$=e^{\epsilon/2}\Pr[S^{(i)}_{\epsilon}(D^{\prime},x_{i})\in T],$ (17)
-which implies that $\text{cost}(S^{(i)}_{\epsilon})-\text{cost}(S_{\epsilon})$ is $\epsilon$-$x_{i}$-exclusive DP, since $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon/2$-$x_{i}$-exclusive DP).
+The output $c$ of the DimReverse algorithm is $\epsilon/2$-differentially private with respect to the input $(X_{1},X_{2},\ldots,X_{k})$, where $X_{i}$ represents the data points in cluster $i$.
+The overall process achieves $\epsilon$-differential privacy through composition, as $(X_{1},X_{2},\ldots,X_{k})$ is partially derived from $Y$, which itself is $\epsilon/2$-differentially private.
+For each explanations $S^{(i)}_{\epsilon}$, let $X,X^{\prime}:X\setminus\{x_{i}\}\sim X^{\prime}\setminus\{x_{i}\}$, i.e., $X$ and $X^{\prime}$ are any two neighbor datasets that differ at exact one data point that is not agent $i$.
+Let $S^{(i)}_{\epsilon}(Y^{X})$ be the value of $S^{(i)}_{\epsilon}$ with input dataset $X$ (and $Y^{X}$ as the private coreset of $X$ respectively).
+Note that we specify the dataset $X$ (and $X^{\prime}$) as a parameter of $S^{(i)}_{\epsilon}$ to highlight the original dataset where the explanation comes from (either $X$ or $X^{\prime}$).
+Fix any set $S$, let $T=\{Y:S^{(i)}_{\epsilon}(Y)\in S\}$ ,i.e., the set of coresets $Y$ that make $S^{(i)}_{\epsilon}(Y)\in S$.
+Since $X\sim X^{\prime}$, we have:
+$\Pr[S^{(i)}_{\epsilon}(Y^{X})\in S]$ $=\Pr[Y^{X}\in T]$ (7)
+$\leq e^{\epsilon/2}\Pr[Y^{X^{\prime}}\in T]$ (8)
+$=e^{\epsilon/2}\Pr[S^{(i)}_{\epsilon}(Y^{X^{\prime}})\in S],$ (9)
+which implies that $\text{cost}(S^{(i)}_{\epsilon})-\text{cost}(S_{\epsilon})$ is $\epsilon$-$x_{i}$-exclusion DP, since $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon/2$-$x_{i}$-exclusion DP).
 ∎
-Lemma 7.
-(Full version of Lemma 1) With probability at least $1-\beta$, $(c_{1},\ldots,c_{k})$ is a $(w,t)$-approximation of $OPT$, where:
-$w$ $=w^{\prime\prime}(1+\alpha),$ (18)
-$t$ $=\mathcal{O}_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right)$ (19)
-Proof.
-This result derives from Theorem 48 of Ghazi et al. [2020] (From Eq(20) to Eq(24) of Ghazi et al. [2020]). ∎
-Lemma 8.
-(Full version of Lemma 2) Fix an $i$.
-If $OPT_{i}\geq w^{\prime\prime}(1+\alpha)OPT+t^{(i)}$, then with probability at least $1-2\beta$, cost($S_{\epsilon}$) and cost($S^{(i)}_{\epsilon}$) released by Algorithm  3 and  4 satisfies that $cost(S^{(i)}_{\epsilon})>cost(S_{\epsilon}$)
+See 2
 Proof.
@@
 $=cost(S_{\epsilon})+t^{(i)}-\Omega_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right).$ (13)
+Set $t^{(i)}=\Omega_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right)$ and the Lemma follows. ∎
+Definition 9.
+For $p\geq 1,\alpha>0$, $\lambda_{p,\alpha/2}\overset{def}{=}\frac{1+\alpha/2}{((1+\alpha/2)^{1/p}-1)^{p}}$.
 Lemma 6.
@@
 Let $\mathcal{S}$ denote a random $d^{\prime}$-dimensional subspace of $\mathbb{R}^{d}$ and let $\Pi_{S}$ denote the projection from $\mathbb{R}^{d}$ to $\mathcal{S}$.
-Then, for any $\zeta\in(0,1)$ we have
-$\Pr\left[\|v\|_{2}\approx_{1+\zeta}\sqrt{d/d^{\prime}}\|\Pi_{\mathcal{S}}(v)\|_{2}\right]\geq 1-2\exp\left(-\frac{d^{\prime}\zeta^{2}}{100}\right)$ (24)
+Then, for any $\tau\in(0,1)$ we have
+$\Pr\left[\|v\|_{2}\approx_{1+\tau}\sqrt{d/d^{\prime}}\|\Pi_{\mathcal{S}}(v)\|_{2}\right]\geq 1-2\exp\left(-\frac{d^{\prime}\tau^{2}}{100}\right)$ (14)
 Lemma 7.
@@
 where $\Pi_{\mathcal{S}}(\mathcal{X})$ denotes the partition $(\Pi_{\mathcal{S}}(X_{1}),\ldots,\Pi_{\mathcal{S}}(X_{k}))$.
-Theorem 6.
-(Full version of Theorem 2) Fix an $i$. With probability at least $1-\beta$, cost($S^{(i)}_{\epsilon}$) released by Algorithm 4 is a $(w,t)$-approximation of the optimal cost $OPT_{i}$, in which:
-$w$ $=w^{\prime}(1+\alpha)$
-$t$ $=w^{\prime}\mathcal{O}_{k,p}\left(\frac{2^{O_{p,\alpha}(d)}k^{2}\log^{2}{n}}{\epsilon}.\text{polylog}(\frac{n}{\beta}\right),$
-and $OPT_{i}$ is the optimal $(k,p)$-clustering cost with a center is fixed at position $x_{i}$.
+See 3
 Proof.
-Let $\tilde{\mathbf{X}}=(\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{n})$ and $X=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{n})$, setting $\alpha^{\prime}=0.1\alpha$, applying Lemma 10, we have:
+Let $\tilde{\mathbf{X}}=(\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{n})$, i.e., the projected data after applying the transformation $\Pi_{\mathcal{S}}$.
+Let $X=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{n})$, i.e., $\tilde{\mathbf{X}}$ after being clipped by $1/\Lambda$.
+By setting $\alpha^{\prime}=0.1\alpha$, and applying Lemma 7, we have:
 ${OPT}_{i}^{\tilde{d}}\leq\left(\frac{d^{\prime}}{d}\right)^{p/2}(1+0.1\alpha)OPT_{i}^{d}.$ (16)
@@
 Using Lemma 6, we have:
-$\Pr\left[\|x\|>\frac{1}{1+\zeta}\sqrt{d/d^{\prime}}\|x^{\prime}\|\right]$ $\geq 1-2\exp(-d^{\prime}\zeta^{2}/100).$ (27)
+$\Pr\left[\|x\|>\frac{1}{1+\tau}\sqrt{d/d^{\prime}}\|x^{\prime}\|\right]$ $\geq 1-2\exp(-d^{\prime}\tau^{2}/100).$ (17)
 Since $x$ is in the unit ball, $\|x\|<1$, which leads to:
-$\Pr\left[\|x^{\prime}\|<(1+\zeta)\sqrt{d^{\prime}/d}\right]$ $\geq 1-2\exp(-d^{\prime}\zeta^{2}/100).$ (28)
-Setting $\zeta=\sqrt{\frac{log(n/\beta)}{0.01}}-1$, $\Lambda=\frac{1}{1+\zeta}\sqrt{d/d^{\prime}}=\sqrt{\frac{0.01}{\log{(n/\beta)}}.\frac{d}{d^{\prime}}}$, we have:
-$\Pr[\|x^{\prime}\|<1/\Lambda]$ $\geq 1-2\exp(-\frac{d^{\prime}\zeta^{2}}{100})$ (29)
+$\Pr\left[\|x^{\prime}\|<(1+\tau)\sqrt{d^{\prime}/d}\right]$ $\geq 1-2\exp(-d^{\prime}\tau^{2}/100).$ (18)
+Setting $\tau=\sqrt{\frac{log(n/\beta)}{0.01}}-1$, $\Lambda=\frac{1}{1+\zeta}\sqrt{d/d^{\prime}}=\sqrt{\frac{0.01}{\log{(n/\beta)}}.\frac{d}{d^{\prime}}}$, we have:
+$\Pr[\|x^{\prime}\|<1/\Lambda]$ $\geq 1-2\exp(-\frac{d^{\prime}\tau^{2}}{100})$ (19)
 $>1-2\exp(-d^{\prime}\log(n/\beta))$ (20)
@@
 By union bound on all $i$, then with probability at least $1-2\beta$, $x^{\prime}_{i}=\Lambda\tilde{x}_{i}$ for all $i$.
-Since $Y$ is the output of PrivateCoreset with input $X$, then by Theorem 38 of Ghazi et al.
-[2020], $Y$ is a $(0.1\alpha,t)$-coreset of $X^{\prime}$ (with probability at least $1-\beta$), with $t^{\prime}$ as:
-$t^{\prime}=O_{p,\alpha}\left(\frac{2^{O_{p,\alpha}(d^{\prime})}k^{2}\log^{2}{n}}{\epsilon}\log\left(\frac{n}{\beta}\right)+1\right)$ (32)
+Since $Y$ is the output of PrivateCoreset with input $X^{\prime}$ and $\zeta$, then by Theorem 38 of Ghazi et al.
+[2020], $Y$ is a $(0.1\alpha,t)$-coreset of $X^{\prime}$ (with probability at least $1-\beta$), with $\alpha=\frac{(100\zeta)^{p}}{10\lambda_{p,\alpha/2}}$ and $t^{\prime}$ as:
+$t^{\prime}=O_{p,\alpha}\left(\frac{2^{O_{p,\alpha}(d^{\prime})}k^{2}\log^{2}{n}}{\epsilon}\log\left(\frac{n}{\beta}\right)+1\right).$ (22)
+We note that alternatively, given a target approximation parameter $\alpha$, we can set $\zeta=0.01\left(\frac{\alpha}{10\lambda_{p,\alpha/2}}\right)^{1/p}$.
 Let $(y_{1},y_{2},\ldots,y_{k})$ be the solution of NonPrivateApproxFC in peCluster for a fixed $i$, $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ be the optimal solution of the clustering with fixed center at $x^{\prime}_{i}$ on $X$, $OPT_{Y}$ be the optimal cost of the clustering with fixed center at $x^{\prime}_{i}$ on $Y$.
@@
 $=w^{\prime}(1+\alpha)OPT^{d}_{i}+w^{\prime}\Theta\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right),$ (36)
-where in $(a)$ we substitute the value of $t^{\prime}$, and $(b)$ is because $d^{\prime}=O_{\alpha}(p^{4}\log(n/\beta))$, and $(c)$ is because $\beta<1$, and the Lemma follows. ∎
-Set $t^{(i)}=\Omega_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right)$ and the Lemma follows. ∎
+where in $(a)$ we substitute the value of $t^{\prime}$, and $(b)$ is because $d^{\prime}=O_{\alpha}(p^{4}\log(n/\beta))$, and $(c)$ is because $\beta<1$, and the Lemma follows.
 Appendix C Additional proofs for $k$-median algorithm
 Definition 10.
-The solution of the $k$-median problem (with demands and a center fixed at a location $C$) can be formulated as finding the optimal solution of the following Integer program (IP):
+The solution of the $k$-median problem (with demands and a center fixed at a location $z$ can be formulated as finding the optimal solution of the following Integer program (IP):
 minimize $\sum_{i,j\in N}d_{j}c_{ij}x_{ij}$ (38)
@@
 $y_{i}$$\in\{0,1\}\text{ for each $i\in N$}$ (43)
-$y_{C}$ $=1\text{ for a fixed $C\in N$}$ (54)
-$x_{CC}$ $=1\text{ for a fixed $C\in N$}.$ (55)
+$y_{z}$ $=1\text{ for a fixed }z\in N$ (44)
+$x_{zz}$ $=1\text{ for a fixed }z\in N.$ (45)
 Lemma 8.
@@
 ∎
-Lemma 12.
-(Lemma 4 of Charikar et al.
-[1999]) For any feasible integer solution $(x^{\prime},y^{\prime})$ for the input with modified demands, there is a feasible integer solution for the original input with cost at most $4\bar{C}_{LP}$ plus the cost of $(x^{\prime},y^{\prime})$ with demands $d^{\prime}$.
-Lemma 13.
-(Theorem 6 of Charikar et al. [1999]) There is a $1/2$-restricted solution $(x^{\prime},y^{\prime})$ of cost at most $2\bar{C}_{LP}$.
 Appendix D Additional proofs for $k$-means algorithm
@@
 Lemma 10.
-(Full version of Lemma 5) Let $S$ be $1$-stable set and $O$ be the optimal set of $k$ centers, we have $\Delta(O)-3\Delta(S)+2R\geq 0$, where $R=\sum_{q\in P}\Delta(q,s_{o_{q}})$.
+Let $S$ be $1$-stable set and $O$ be the optimal set of $k$ centers, we have $\Delta(O)-3\Delta(S)+2R\geq 0$, where $R=\sum_{q\in P}\Delta(q,s_{o_{q}})$.
 Proof.
@@
 Lemma 12.
-(Full version of Lemma 6) With $R$ and $\alpha$ defined as above: $R\leq 2\Delta(O)+(1+2/\alpha)\Delta(S)$.
+With $R$ and $\alpha$ defined as above: $R\leq 2\Delta(O)+(1+2/\alpha)\Delta(S)$.
 Proof.
@@
 ∎
-Figure 2: A detailed visualization of our dataset (Charlottesville County, Virginia) and analysis includes (a) a Scatter plot of the full dataset with sampled points for contrastive analysis.
+D.3 Additional details on experiments
+D.3.1 Datasets and Experimental Setup
+Figure 3: A detailed visualization of our dataset (Charlottesville County, Virginia) and analysis includes (a) A scatter plot of the full dataset with 100 randomly selected points for contrastive analysis, chosen to provide a more comfortable and manageable subset for explanation purposes.
 (b) Comparison of $k$-median clustering with fixed and non-fixed centroids, both private and non-private.
 (c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-median with a fixed centroid.
-(d) Heatmap indicating $k$-median error costs based on fixed centroid choices.
+Figure 4: The figure presents visualizations of the dataset (Heart Disease, UCI MLR) reduced to four dimensions (2, 4, 6, and 8), along with the corresponding analysis (a) A t-SNE scatter plot illustrating the high-dimensional data.
+(b) Comparison of $k$-means clustering with fixed and non-fixed centroids, both private and non-private.
+(c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-means with a fixed centroid.
+Figure 5: The figure presents further visualizations of the dataset (Heart Disease, UCI MLR) introduced in the main paper, reduced to four dimensions (2, 4, 6, and 8), along with the corresponding analysis (a) A t-SNE scatter plot illustrating the high-dimensional data.
+(b) Comparison of $k$-median clustering with fixed and non-fixed centroids, both private and non-private.
+(c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-median with a fixed centroid.
+Figure 6: The figure presents visualizations of the dataset (Breast Cancer - 30 features, UCI MLR) reduced to four dimensions (4, 6, 8, and 16), along with the corresponding analysis (a) A t-SNE scatter plot illustrating the high-dimensional data.
+(b) Comparison of $k$-means clustering with fixed and non-fixed centroids, both private and non-private.
+(c) Bar graph showing contrastive explanation differences for differential private and non-private $k$-means with a fixed centroid.
+In the main body of the paper, we focused on the Heart Disease dataset from the UCI Machine Learning Repository, which was reduced from 13 features to 8 features for visualization purposes.
+However, we have also conducted experiments on the same dataset with different dimensionality reductions to analyze the impact on our results.
+Furthermore, to demonstrate the robustness of our work, we have applied our analysis to additional datasets, including synthetic datasets and real-world datasets from Charlottesville, Albemarle and Breast Cancer Wisconsin dataset including 30 features.
+This appendix provides a detailed discussion of these experiments and their outcomes.
+Real-World Datasets
+To demonstrate the robustness and real-world applicability of our approach, we conducted experiments on real-world datasets: the Charlottesville City dataset, the Albemarle County dataset, Heart Disease and Breast Cancer Wisconsin.
+Charlottesville City Dataset
+The Charlottesville City dataset is part of the synthetic U.S.
+population, as described in [Chen et al., 2021] and [Barrett et al., 2009].
+This dataset consists of approximately 33,000 individuals and around 5,600 activity locations visited by these individuals.
+The locations represent places where individuals perform various activities, providing insights into human mobility patterns and social interactions within the city.
+Albemarle County Dataset
+The Albemarle County dataset is another real-world dataset used in our experiments.
+This dataset is significantly larger than the Charlottesville City dataset, comprising about 74,000 individuals.
+The increased sample size allows us to evaluate the scalability and performance of our approach when applied to larger, more complex datasets.
+The Albemarle County dataset contains information about individuals’ activities and the locations they visit, similar to the Charlottesville City dataset.
+This dataset provides a comprehensive representation of human mobility patterns and social interactions within the county.
+By using these real-world datasets, we aim to validate the effectiveness and practicality of our methodology in real-life scenarios.
+The diverse characteristics of these datasets, such as the number of individuals and activity locations, enable us to assess the robustness and generalizability of our approach.
+In our experiments, we applied our methodology to all datasets and compared the results to those obtained from the synthetic datasets.
+The consistency of results across these real-world datasets further reinforces the reliability and potential of our approach for real-world applications.
+Synthetic 2D Dataset:
+We carefully created a synthetic dataset that mimics the properties of real-world datasets, striking a balance between realism and controlled variability.
+This dataset consists of 1000 uniformly distributed data points in a 2D space, with a range similar to the real datasets we analyzed.
+The primary motivation behind this synthetic dataset is to provide a sandbox environment free from the unpredictable noise and anomalies of real-world data.
+This controlled setting is pivotal in understanding the core effects of differential privacy mechanisms, isolating them from external confounding factors.
+The dataset is a foundational tool in our experiments, allowing us to draw comparisons and validate our methodologies before applying them to more complex, real-world scenarios.
 Figure 7: A detailed visualization of our synthetic dataset, measured by the same metrics as the other real-world datasets.
-D.3 Additional details on experiments
-D.3.1 Datasets and Experimental Setup
-Beyond the Albemarle County dataset discussed in the main body of the paper, we also conducted experiments on other datasets, including the Charlottesville city dataset and synthetic activity-based populations.
-Here, we provide a detailed analysis of these experiments.
-Charlottesville City Dataset:
-This dataset is part of the synthetic U.S.
-population, as described in  Chen et al.
-[2021], Barrett et al.
-[2009].
-It comprises about 33K individuals and approximately 5600 activity locations visited by these individuals.
-The locations represent places where individuals perform various activities.
-Synthetic 2D Dataset:
-The synthetic dataset was meticulously crafted to emulate the distribution and characteristics of real-world datasets, ensuring a balance between realism and controlled variability.
-Comprising 100 data points in a 2D space, this dataset was uniformly distributed, mirroring the range observed in the real datasets we analyzed.
-The primary motivation behind this synthetic dataset is to provide a sandbox environment, free from the unpredictable noise and anomalies of real-world data.
-This controlled setting is pivotal in understanding the core effects of differential privacy mechanisms, isolating them from external confounding factors.
-The dataset serves as a foundational tool in our experiments, allowing us to draw comparisons and validate our methodologies before applying them to more complex, real-world scenarios.
 D.3.2 Experimental Results
 Impact of $\epsilon$ on Private Optimal (PO) and Private Contrastive (PC):
-For the Charlottesville city dataset, we observed similar trends to those in the Albemarle county dataset.
-As the $\epsilon$ value increased, prioritizing accuracy, there was a slight compromise in privacy.
-However, consistent with our hypothesis, the influence of the epsilon budget on the explainability of our outcomes was negligible.
-Effect of Privacy on Explainability:
-The contrastive explanations for the Charlottesville city dataset also showcased the resilience of our approach.
-Even with differential privacy perturbations, the quality of explanations remained consistent, emphasizing the robustness of our method.
-Visualization:
-Figure  2 - (a) showcases the Charlottesville city dataset.
-Blue dots represent the original dataset, while orange ’X’ markers indicate the sampled points for contrastives.
-The distribution of these points provides a visual representation of the data’s diversity and the regions we focused on for our contrastive analysis.
+We observed consistent trends and patterns in all our data sets, including Charlottesville city, Albemarle county, and notably the Heart Disease dataset and the Breast Cancer dataset from the UCI Machine Learning Repository.
+As the value of $\epsilon$ increased to prioritize accuracy, we observed a gradual reduction in privacy protection.
+However, in line with our hypothesis, the impact of the epsilon budget on the explainability of our outcomes remained minimal.
+This consistency held true across all dimensions of the Heart Disease dataset, reinforcing the robustness of our findings across diverse data sources and attributes.
 D.3.3 Performance Evaluation:
-For each $\epsilon$ value, we conducted 25 different runs on the Charlottesville city dataset.
-The average results were consistent with our findings from the Albemarle County dataset.
+For each $\epsilon$ value, we conducted 100 different runs for each dataset.
+The average results were consistent with our findings across all datasets.
 It’s essential to note that these multiple invocations were solely for performance evaluation.
@@
 Consistency in Contrastive Explanations across Datasets:
-Despite the distinct scales between the Charlottesville city and the Albemarle county datasets, we observed consistent patterns in the contrastive explanations.
-Specifically, as illustrated in Figure 2 - (b) and Figure  2 - (c), the contrastive explanations remained largely unaffected by variations in the $\epsilon$ budget.
+Despite the distinct scales between our different datasets, we observed consistent patterns in the contrastive explanations.
+Specifically, as illustrated in all Figures - (b), contrastive explanations remained largely unaffected by variations in the $\epsilon$ budget.
 This consistency further reinforces our hypothesis that the epsilon budget has a negligible influence on the explainability of our outcomes, even when applied to datasets of different scales.
-Impact of Fixed Centroids on Contrastive Error:
-In Figure  2 - (d), we delve deeper into the effects of selecting different centroids as fixed centroids.
-It’s evident that the choice of a fixed centroid has a pronounced impact on the contrastive error.
-This underscores the importance of the centroid’s position concerning the data distribution.
-Some centroids, especially those closer to data-dense regions, tend to produce lower counterfactual errors, while others, particularly those near sparser regions or outliers, result in higher errors.
-This observation highlights the intricate relationship between centroid positioning and the resulting contrastive explanation’s accuracy.
 D.3.4 Conclusion:
-The extended experiments on the Charlottesville city dataset further validate our approach’s efficacy.
+The extended experiments on all our datasets further validate our approach’s efficacy.
 The balance between privacy and utility, the robustness of contrastive explanations, and the negligible impact of $\epsilon$ on explainability were consistent across datasets.